### v3

# Did I Faithfully Say What I Thought? Bridging the Gap Between LLMs Neural Activity and Self-Explanations

Abstract
Large Language Models (LLMs) can generate plausible free text self-explanations to justify their answers. However, these natural language explanations may not accurately reflect the model’s actual reasoning process, indicating a lack of faithfulness. Existing faithfulness evaluation methods rely primarily on behavioral tests or computational block analysis without examining the semantic content of internal neural representations. This paper proposes NeuroFaith, a flexible framework that measures the faithfulness of LLM free text self-explanation by identifying key concepts within explanations and mechanistically testing whether these concepts actually influence the model’s predictions. We show the versatility of NeuroFaith across 2-hop reasoning and classification tasks. Additionally, a linear faithfulness probe based on NeuroFaith is developed to detect unfaithful self-explanations from representation space and improve faithfulness through steering. NeuroFaith provides a principled approach to evaluating and enhancing the faithfulness of LLM free text self-explanations, addressing critical needs for trustworthy AI systems.
1 Introduction
Autoregressive Large Language Models (LLMs) can generate plausible self Natural Language Explanation (self-NLE) to support their answers (Wiegreffe & Marasovic, ; Huang et al., 2023). Generating self-NLE consists of prompting the LLM to generate an explanation in a predict-then-explain setting, where the model first generates a response to a question and then produces a self-NLE as a justification. Unlike their non-generative predecessors, modern LLMs are trained to generate both answers and free text self-NLE that appear credible despite potentially containing persuasive hallucinations (Sahoo et al., 2024). This way, despite their logical and coherent appearance favoring trust in the model (Han et al., 2023), LLM-generated self-NLE turn out to not systematically reflect the actual underlying decision-making process of the model, creating a tension between self-NLE plausibility and faithfulness (Agarwal et al., 2024).
Faithfulness, as defined by Jacovi & Goldberg (2020), measures ”how accurately the explanation reflects the true reasoning process of the model”, a definition widely adopted in the literature (Lyu et al., 2024) and which we likewise follow throughout this work. Unfaithful self-NLE can have serious consequences in critical domains, where explanations that appear plausible but lack faithfulness might lead end-users to over-rely on model predictions and make unfair (Luo et al., 2022) or harmful (Kayser et al., 2024) decisions. As the use of LLMs is expanding across diverse fields, the combination of their widespread adoption and the simplicity of generating self-NLE through prompting make evaluating their faithfulness increasingly critical.
Assessing self-NLE faithfulness presents profound difficulties. Unlike more structured explainability methods such as attribution (Wiegreffe et al., 2021) or counterfactual approaches (Madsen et al., 2024), the free-form nature of natural language explanations makes faithfulness evaluation particularly difficult. Numerous methods have been developed to measure self-explanation faithfulness (see e.g. Lyu et al. (2024)). However, (1) they mostly perform behavioral tests and do not examine LLM internal reasoning processes (Atanasova et al., 2023; Siegel et al., 2024, 2025; Matton et al., 2025), and (2) they identify the computational blocks that contribute to prediction and self-NLE without conducting semantic analysis of the neural representations within these blocks (Wiegreffe et al., 2021; Parcalabescu & Frank, 2024; Yeo et al., 2024). These shortcomings have led these methods to be characterized as measuring self-consistency rather than genuine faithfulness (Parcalabescu & Frank, 2024), since they fail to establish direct connections between explanations and the model’s actual reasoning processes. To overcome these limitations, we introduce NeuroFaith, a flexible framework for directly measuring LLM self-explanation faithfulness. Our main contributions are as follows:
  1. 1.
    NeuroFaith measures the alignment between LLM internal reasoning and its self-explanations by identifying key concepts within the explanation and mechanistically testing whether these concepts actually influence the model’s predictions.
  2. 2.
    We demonstrate the versatility of NeuroFaith by applying it to both 2-hop reasoning tasks and classification problems.
  3. 3.
    We develop a linear faithfulness probe based on NeuroFaith to efficiently detect unfaithful LLM self-explanations from representation space and improve faithfulness through steering.
This paper is organized as follows: Section 2 presents how existing approaches measure the faithfulness of LLM self-NLE. Section 3 introduce the NeuroFaith framework. In Sections 4 and 5, we instantiate NeuroFaith for two different tasks, 2-hop reasoning and classification respectively. We finally show in Section 6 that self-NLE faithfulness, as measured by NeuroFaith, can be linearly detected in LLM representation space and improved through steering.
2 Related Work
Numerous approaches have been proposed to measure self-NLE faithfulness (Lyu et al., 2024). One approach (Tutek et al., 2025) assesses the effect of unlearning (Liu et al., 2025) the parametric knowledge encoded in the reasoning steps of the self-NLE. The higher the change in prediction between the original model and the model having unlearned the reasoning steps, the more faithful the self-NLE. We group the remaining approaches in two categories.
Counterfactual Interventions.
NLE faithfulness can be assessed through behavioral tests that measure how perturbations in the input text affect both predictions and self-NLE (Atanasova et al., 2023). Counterfactual Intervention (CI) methods employ auxiliary models to generate counterfactual texts designed to change the LLM outcome. The LLM is then prompted to produce a self-NLE to justify its new prediction. The self-NLE is deemed faithful if it aligns with the specific CI that caused the prediction change. These CI approaches mostly differ in two ways: how they measure consistency between the intervention and the resulting self-NLE (Atanasova et al., 2023; Siegel et al., 2024, 2025), and the granularity of the CI intervention (Matton et al., 2025). These approaches face several limitations: (1) the CI may not be solely responsible for the change in prediction, as the model might base its new prediction on another part of the input text after intervention, (2) CI methods treat the model as a black box by only examining input-output relationships without analyzing the internal neural processes that generate predictions, which departs from the commonly adopted definition of explanation faithfulness.
Attribution Agreement.
Another way to measure self-NLE faithfulness is to compute post-hoc Attribution Agreement (AA) between the prediction and the NLE (Parcalabescu & Frank, 2024; Wiegreffe et al., 2021; Yeo et al., 2024). AA methods compute attribution scores for both the model’s predictions and its self-NLE and then measure the correlation between these scores to assess faithfulness. Higher correlation values indicate greater faithfulness in the model’s self-NLE. AA approaches vary in the post-hoc attribution method employed (gradient-based (Sundararajan et al., 2017), SHAP (Lundberg & Lee, 2017) or activation patching (Meng et al., 2022)). While AA methods assess whether the same LLM computational blocks were used when generating both the prediction and the self-NLE, they overlook the semantic content of neural representations, leaving the model’s reasoning process only partially treated.
In the following, we propose a framework that directly examines the correspondence between self-NLE and the model’s actual reasoning process by conducting concept-level mechanistic analysis of internal hidden states during the forward pass that generates the prediction.
3 NeuroFaith: A Framework for Measuring the Faithfulness of LLM self-NLE
This section introduces the core principles of NeuroFaith, our proposed flexible framework for measuring the faithfulness of LLM self-natural language explanations (self-NLE). Based on the premise that faithful explanations should accurately reflect the model’s internal reasoning process, NeuroFaith quantifies how well a self-NLE aligns with the model’s internal reasoning processes by extracting concepts from self-NLE and mechanistically evaluating their importance for the model prediction. Sections 4 and 5 provide detailed instantiations of how to apply NeuroFaith to 2-hop reasoning and classification tasks respectively. In the following, we denote an $L$-layer auto-regressive Transformer-based LLM $f$, a set of input texts X and a text of interest $x\in\textbf{X}$. We denote $f(x)$ as the model’s answer and $e(x)$ as the corresponding self-NLE produced by $f$ for input $x$. NeuroFaith is a 3-step framework illustrated in Figure 1, summarized below and detailed in the next subsections.
  1. 1.
    Concept Extraction. We extract from $e(x)$ a set of concepts $\left\{c_{i}\right\}_{i=1}^{p}$ that influence $f(x)$.
  2. 2.
    Concept-wise Mechanistic Interpretation. For each concept $c_{i}$, we generate a post-hoc interpretation $\texttt{I}_{\Gamma}(c_{i})$ based on a relevant subpart of the model called circuit $\Gamma$ and a mechanistic interpretability method called interpreter I to assess its impact on $f(x)$.
  3. 3.
    Faithfulness Measurement. We compute faithfulness $F(x,e)$ by validating that the concepts $\left\{c_{i}\right\}_{i=1}^{p}$ emphasized in the self-NLE actually have significant mechanistic effects on $f(x)$ as determined by their interpretations $\left\{\texttt{I}_{\Gamma}(c_{i})\right\}_{i=1}^{p}$.
3.1 Concept Extraction
Given an input text $x$, a prediction $f(x)$ and its self-NLE $e(x)$, the first step involves parsing the self-NLE to extract a set of concepts $\left\{c_{i}\right\}_{i=1}^{p}$ that influence $f(x)$.
Concept Definition. Recent research has shown that interpretable binary high-level features, referred to as concepts, appear to be linearly encoded within LLM representation space (Elhage et al., 2022; Park et al., 2024). This computational understanding aligns with definitions from cognitive science (Ruiz Luyten & van der Schaar, 2024), where concepts are considered as mental entities essential to thought, enabling information integration and categorization (Goguen, 2005). This alignment makes concepts an ideal granularity level for model interpretation. For example in Figure 1, the concept ”Ingmar Bergman” directly influences the prediction ”Sweden” because Ingmar Bergman is Swedish and directed the movie Persona.
Concept Extraction. Concepts can be extracted either through human investigation or using an auxiliary LLM under LLM-as-a-Judge settings (Gu et al., 2024). Human annotation enables targeted analysis of specific concepts that are hypothesized to be important for the model’s reasoning. Human annotation provides high-quality, domain-expert identification of relevant influential concepts but is costly and may not scale to large datasets. LLM-as-a-Judge approaches offer scalability but may introduce systematic biases or miss relevant concepts that human experts would identify. We provide detailed examples of prompts used to extract concepts from self-NLE with an auxiliary LLM in Appendix D.1.
Figure 1: NeuroFaith overview. NeuroFaith (1) extracting concepts from the self-NLE and (2) assessing the mechanistic influence of these concepts to finally (3) measure faithfulness.
3.2 Concept-wise Mechanistic Interpretation
The second step evaluates whether the concepts $\left\{c_{i}\right\}_{i=1}^{p}$ are actually important for the model’s internal processing. We compute an attribution score $\texttt{I}_{\Gamma}(c)$ for each concept $c\in\left\{c_{i}\right\}_{i=1}^{p}$ using an interpretability method (interpreter I) applied to model hidden states within a relevant subpart of the model (circuit $\Gamma$). We assume each concept can be detected in the model representation space.
Circuit.
Rather than examining every network component, we focus our mechanistic analysis on computational subgraphs that most significantly influence $f$ prediction named circuits (Elhage et al., 2021). Circuits are sparse oriented subgraphs with interpretable functional roles, where nodes represent computational units and edges represent computation paths (Räuker et al., 2023). Formally, we define a circuit as $\Gamma=(\left\{(k,\ell)\right\},\mathcal{G})$ where $\left\{(k,\ell)\right\}$ specifies coordinate pairs (token index, layer) for information flow in $f$, and $\mathcal{G}$ defines the granularity of each computational node. The granularity $\mathcal{G}$ represents transformer sublayers: residual stream output (RS), multi-head attention (MHA), or multi-layer perceptron (MLP(Rai et al., 2024). This granularity avoids focusing on individual neurons due to their polysemanticity (Elhage et al., 2022). For example in Figure 1, the circuit is highlighted in blue, starting at token index 12 and layer 1 and finishes at token index 13 and layer L-1. Circuits can be obtained through manual investigation of task-specific patterns (Bereska & Gavves, 2024; Wang et al., 2023; Biran et al., 2024) or automated discovery via activation patching (Meng et al., 2022; Conmy et al., 2023), backward attribution (Ferrando & Voita, 2024), or Transcoders (Dunefsky et al., 2024). Additional details about circutis are in Appendix D.2.
Mechanistic Interpreter.
To mechanistically interpret a concept $c$ across circuit $\Gamma$, we use an interpreter I to compute an attribution score $\texttt{I}_{\Gamma}(c)$ that evaluates whether $c$ impacts $f(x)$. NeuroFaith implements two approaches, each suited to different objectives. Probing methods determine whether $c$ is represented within $f$ hidden states on $\Gamma$. This involves either generating natural language interpretations of hidden states (using Selfie (Chen et al., 2024) or Patchscopes (Ghandeharioun et al., 2024)) to detect $c$, or training linear probes (Belinkov, 2022) when $c$ is linearly separable with sufficient labels to train the probe. Denoting $h^{\ell}_{k}$ the hidden state at token $k$ and layer $\ell$, we define $\texttt{I}_{\Gamma}(c)=\max_{(k,\ell)\in\Gamma}\texttt{I}(h^{\ell}_{k})$, assessing $c$ as important if detected in at least one hidden state from $\Gamma$. Concept importance methods approximate the causal influence of $c$ on $f(x)$ (Geiger et al., 2025). They use gradient-based methods like TCAV (Kim et al., 2018) or Representation Engineering (Zou et al., 2023) to directly manipulate concept-related activations across $\Gamma$ and measure behavioral changes, providing causal evidence for concept importance. Section 5 details this procedure for measuring faithfulness in the case of classification.
3.3 Faithfulness Measurement
We consider the self-NLE $e(x)$ as faithful when the extracted concepts $\left\{c_{i}\right\}_{i=1}^{p}$ are demonstrably important for $f$’s internal processing according to $\texttt{I}_{\Gamma}(c)$. We propose to define the faithfulness of $e(x)$ as the proportion of concepts it contains that are mechanistically assessed as important, i.e. that have a positive attribution score: $F(x,e)=\frac{1}{p}\sum_{i=1}^{p}\mathbf{1}_{\texttt{I}_{\Gamma}(c_{i})>0}$.
The faithfulness score has different meanings depending on the chosen interpreter I: Probing-based faithfulness measures the detection rate of explanation concepts within the model’s internal representations along circuit $\Gamma$. High probing-based faithfulness indicate that most mentioned concepts are properly decoded from the model’s hidden states. Importance-based faithfulness estimate causal relevance of explanation concepts for the prediction $f(x)$. High importance-based faithfulness indicate that most mentioned concepts actually influence $f$ reasoning process.This framework directly captures concept-level alignment between the self-NLE $e(x)$ and $f$ internal reasoning process. NeuroFaith faithfulness scores either indicate that the self-NLE accurately reflects internal processing or that the explanation mentions concepts that are not mechanistically important. This way, NeuroFaith directly aligns with the established definition of explanation faithfulness introduced above (Jacovi & Goldberg, 2020), offering a principled evaluation of self-NLE faithfulness.
4 The Case of 2-Hop Reasoning
In the previous section, we defined the high level core principle of NeuroFaith. We now instantiate NeuroFaith to 2-hop reasoning using probing-based faithfulness, and precisely characterize self-NLE with respect to both its faithfulness and its correctness.
4.1 NeuroFaith Instantiation for 2-Hop Reasoning
Task Description.
[…]
Concept Extraction.
[…]
Probing-based Concept Interpretation.
The presence of a single bridge object in 2-hop reasoning self-NLE makes probing-based interpretations particularly appropriate. For $f$ to correctly answer, it must internally compute the bridge object during the first hop before executing the second hop. Therefore, detecting the extracted bridge object $\widehat{o_{2}}=c$ in $f$ internal representations is a strong indication to assess whether $c$ is important for the prediction. We employ natural language interpretation methods such as Selfie and Patchscopes, rather than linear probes, as they provide higher accuracy and are unsupervised (Ghandeharioun et al., 2024). Recent work (Biran et al., 2024; Yang et al., 2024) has shown that the bridge object of 2-hop reasoning was resolved on early layers and can be detected by focusing on the last token corresponding to source object $o_{1}$ and the residual stream (RS). We leverage this information to define circuit $\Gamma$ and generate the natural language description $\tilde{h^{\ell}_{k}}$ of hidden state $h^{\ell}_{k}$. For concept (bridge object) $c$, the probing-based concept attribution is defined as $\texttt{I}_{\Gamma}(c)=1$ iff $\exists(k,\ell)\in\Gamma$ such that $c\in\tilde{h^{\ell}_{k}}$.
Self-NLE and Latent Reasoning Characterization.
Following the notations introduced in Section 3.3, $e(x)$ is assessed faithful if $c$ is detected in at least one hidden state within $\Gamma$ with the probe I ($\texttt{I}_{\Gamma}(c)$=1). Beyond faithfulness, we can use the ground truth object ($o_{2}$) to characterize the correctness of $e(x)$ and the latent first hop respectively: $e(x)$ is said to be correct (self-NLE correctness) if its corresponding bridge object is the expected ground truth ($c=o_{2}$). Likewise, the model performs correct first-hop latent reasoning (latent hop 1 correctness) if $\exists(k,l)\in\Gamma$ such that $o_{2}\in\tilde{h^{\ell}_{k}}$. This multi-dimensional analysis provides comprehensive characterization of explanation quality and internal reasoning alignment. We give a thorough taxonomy of self-NLE in the case of 2-hop reasoning in Appendix E, based on prediction correctness, self-NLE faithfulness and correctness and latent reasoning correctness. We also propose an experimental protocol to evaluate the relevance of our faithfulness measure in Appendix F and compare NeuroFaith to CI, showing our faithfulness measure relevancy in absolute and in comparison to CI.
4.2 2-Hop Reasoning Experimental Analysis
Experimental Setup.
We evaluate NeuroFaith for 2-hop reasoning on the Wikidata-2-hop dataset (Biran et al., 2024). We apply NeuroFaith to Gemma-2-2B, Gemma-2-9B and Gemma-2-27B (Riviere et al., 2024). We use Qwen3-32B as auxiliary model to extract bridge objects from self-NLE and use the prompt from Ghandeharioun et al. (2024) to apply Patchscopes as interpreter. We evaluate Latent Reasoning Hop 1 Correctness, Self-NLE Faithfulness and Self-NLE Correctness based on the characterization introduced above. The results are computed according to the status of the prediction preceding the self-NLE (accurate vs. inaccurate). We give more details about the layers used to define $\Gamma$ and the prompt used to extract the bridge object and compute the interpretations in Appendix D.1 and  D.2.
Model Task Acc. Self-NLE Latent Hop 1 Self-NLE
Correctness Correctness Faithfulness
Accurate Inaccurate Accurate Inaccurate Accurate Inaccurate
gemma-2-2b 7.1% 58.0% 56.0% 48.3% 47.4% 48.4% 57.6%
gemma-2-9b 16.6% 74.0% 55.4% 58.5% 44.3% 60.8% 55.0%
gemma-2-27b 22.3% 77.8% 55.6% 64.9% 46.9% 68.8% 60.2%
Table 1: 2-hop reasoning accuracy, first hop latent reasoning correctness and self-NLE correctness and faithfulness across models obtained from NeuroFaith. ”(In)accurate” represents the set of predictions initially (in)correct.
Key Findings.
Table 1 shows the aggregated experimental results obtained by applying NeuroFaith on Wikidata-2-hop. All the analyzed metrics are higher on average for accurate predictions and improve with model size. These results support that correct reasoning tend to produce more faithful and correct self-NLE. Key failure patterns emerge: approximately 45% of inaccurate predictions correctly resolve the first hop operation, indicating that failure often occurs in the second reasoning step ($o_{2},r_{2},o_{3}$). This suggests that models can identify correct bridge objects even when failing to complete the full reasoning chain. The difference between latent reasoning hop 1 correctness and self-NLE correctness shows that models sometimes identify correct bridge objects internally but fail to express them in their self-NLE. This gap decreases with model size, suggesting better alignment between internal and explicit reasoning in larger models.
Remarkably, even when models produce correct final answers, they identify the correct bridge object in their internal representations only 48-65% of the time. This suggests that models can arrive at correct answers through alternative reasoning pathways that bypass explicit bridge object computation. The models may be leveraging (1) direct associations between source and target entities (shortcut learning (Geirhos et al., 2020)), or (2) alternative pathways making the model accurately answer for a reason different than the ground truth bridge object (see category 9 in Figure 25). This phenomenon highlights that task accuracy alone is insufficient for evaluating reasoning quality.
5 The Case of Classification
This section instantiates NeuroFaith for classification problems by employing concept importance methods as the mechanistic interpreter I. For classification tasks, given an input text $x\in\textbf{X}$, we denote $f(x)=\hat{y}\in\mathcal{Y}$ with probability score $p_{\hat{y}}(x)$ where $\mathcal{Y}$ is the label space. We assume access to a predefined set of task-relevant concepts $\mathcal{C}$ with concept labels for input texts.
5.1 NeuroFaith Instantiation for Classification
Concept Extraction.
To enable mechanistic interpretation of each concept $c\in\mathcal{C}$, we compute Concept Activation Vectors (CAVs) that represent concepts as directions in $f$ representation space. Following established practices in concept-based interpretability, we employ the mean difference (diff-mean(Rimsky et al., 2024) approach for CAV computation due to its optimal balance between concept detection accuracy and computational efficiency (Wu et al., 2025). For a concept $c\in\mathcal{C}$, a token index $k$ and a layer $\ell$, the layer-wise CAV is defined as: $\overrightarrow{c_{\ell}}=\frac{1}{\left|\textbf{X}^{+}_{c}\right|}\sum_{x\in\textbf{X}^{+}_{c}}{h^{\ell}_{k}}-\frac{1}{\left|\textbf{X}^{-}_{c}\right|}\sum_{x\in\textbf{X}^{-}_{c}}{h^{\ell}_{k}}$, where $\textbf{X}^{+}_{c}$ and $\textbf{X}^{-}_{c}$ respectively represent the sets of texts from X where the concept $c$ is present or absent and $h^{\ell}_{k}$ denotes the hidden state at layer $\ell$ and token position $k$ at the specified granularity $\mathcal{G}$. We set the token index to the final position of $x$, as this location represents the model’s complete computational state prior to next-token generation. These CAVs allow for computing the importance of each concept using Representation Engineering techniques. Following Section 3, we use an auxiliary LLM to extract a set of concepts from $e(x)$, making sure that $\left\{c_{i}\right\}_{i=1}^{p}\subset\mathcal{C}$.
Importance-based Concept Interpretation.
For a concept $c\in\left\{c_{i}\right\}_{i=1}^{p}$, we approximate its causal influence on $f(x)$ through representation engineering and concept erasure (Belrose et al., 2023). We perform controlled interventions by erasing $c$ from hidden states during forward propagation: $h^{\ell}_{k}\leftarrow h^{\ell}_{k}-\lambda\times\overrightarrow{c_{\ell}}$, where $\lambda\in[0,1]$ represents intervention intensity and with $\lambda=1$ represents maximum intervention intensity to avoid $f$ collapse (Rimsky et al., 2024). This approach provides strong causal evidence for concept importance without requiring computationally costly gradient computations such as TCAV. We define $\Gamma$ with granularity $\mathcal{G}$ set to residual stream (RS), token index as the final position (likewise CAVs), and layers are selected based on concept detectability (F1 score >> 60% using layer-wise linear probes with diff-mean). Applying this intervention across circuit $\Gamma$, we measure the resulting probability score: $p_{\hat{y}}(x,\left\{\text{do}(H^{\ell}_{k}=h^{\ell}_{k}-\lambda\times\overrightarrow{c_{\ell}})\right\}_{(k,\ell)\in\Gamma})$, where the $\text{do}(X=x)$ operator (Pearl, 2009) represents an intervention that sets variable $X$ to value $x$. We model this relationship as linear: $p_{\hat{y}}(\cdot)=\beta_{0}+\beta_{1}\times\lambda$, where the concept importance score is $\texttt{I}_{\Gamma}(c)=\beta_{1}$ (i.e. the marginal effect of intervention on prediction probability, computed via linear regression across different $\lambda$ values). $\texttt{I}_{\Gamma}(c)$ is set to 0 when its significance $t$-test shows $p>0.01$. This process is repeated for each extracted concept from $e(x)$ to derive faithfulness scores as described in Section 3.3. Implementation details and examples are provided in Appendix D.2 and G.
Model Task Acc. Self-NLE Faithfulness
AGNews Ledgar AGNews Ledgar
Accurate Inaccurate Accurate Inaccurate
gemma-2-2B 86.5% 40.3% 54.6% 65.1% 58.9% 40.5%
gemma-2-9B 88.7% 59.7% 96.0% 92.3% 56.0% 58.3%
gemma-2-27B 88.9% 58.1% 74.0% 72.4% 53.0% 31.6%
Table 2: Classification accuracy and self-NLE faithfulness across models and datasets obtained from NeuroFaith. ”(In)accurate represents the set of predictions initially (in)correct.
5.2 Classification Experimental Analysis
Experimental Setup.
We evaluate NeuroFaith’s classification instantiation on two datasets with varying complexity: AGNews (Gulli, 2005), a newspaper article classification and Ledgar (Tuggener et al., 2020), a more challenging critical domain legal document classification dataset. We apply NeuroFaith to Gemma-2-2B, Gemma-2-9B and Gemma-2-27B (Riviere et al., 2024). We use the concept set $\mathcal{C}$ and labels of AGNews and Ledgar from Bhan et al. (2025) to compute the CAVs. We use Qwen3-32B to extract concepts from the self-NLE, ensuring extracted concepts belong to $\mathcal{C}$ for each dataset. We focus on instances solely containing concepts that are linearly detectable in at least one layer. We evaluate Self-NLE Faithfulness according to the status of the prediction preceding the self-NLE (accurate vs. inaccurate).
Key Findings.
Table 2 shows experimental results for AGNews and Ledgar datasets. Self-NLE faithfulness is consistently higher on the simpler AGNews dataset, aligning with the intuition that explanation faithfulness might decrease with task complexity. Accurate predictions generally yield more faithful explanations, suggesting that correct predictions might rely on relevant concepts. Notably, gemma-2-9B achieves the highest faithfulness scores across both datasets, indicating better alignment and going against the idea of a monotonic relationship between model scale and explanation faithfulness.
Comparing findings across classification and 2-hop reasoning reveals both general patterns and task-specific differences. General findings: (1) accurate predictions tend to produce more faithful explanations, and (2) increased task difficulty seems to reduce self-NLE faithfulness. However, scaling effects vary between tasks, indicating that the model scale-faithfulness relationship is highly task-dependent, a finding consistent with prior work (Madsen et al., 2024; Atanasova et al., 2023; Parcalabescu & Frank, 2024; Matton et al., 2025).
6 Linearly Detecting and Improving Self-NLE Faithfulness
Recent work has demonstrated that various AI safety behaviors naturally associated with faithfulness are encoded as linear directions in LLM representation spaces (Bereska & Gavves, 2024), such as hallucination (Rimsky et al., 2024) or deceptiveness (Goldowsky-Dill et al., 2025). We investigate whether NeuroFaith-based faithfulness exhibits linear structure enabling (1) accurate detection in representation space and (2) manipulation to improve faithfulness through activation steering.
6.1 Linear Detection of Faithfulness
Figure 2: Majority vote faithfulness linear probe performance across models and datasets.
Faithfulness Linear Representation Computation.
We construct datasets of faithful and unfaithful self-NLE pairs using NeuroFaith scores. Unlike 2-hop reasoning tasks, classification self-NLE faithfulness yields continuous scores between 0 and 1. To ensure clear class separation, we focus on polarized cases where $F(x)\in\left\{0,1\right\}$ for clear separation. We construct input sequences as $x_{nle}=[x,\hat{y},e(x)]$ and extract hidden states from final token positions across all layers. We train diff-mean linear probes on these states, yielding layer-wise faithfulness representations $\overrightarrow{F_{\ell}}$. For classification tasks, we use class-averaged faithfulness vectors to isolate faithfulness-specific representations independent of class confounders. Based on these faithfulness vectors, we compute a majority vote across layer-wise linear probes beyond layer 5. Figure 2 shows faithfulness can be reliably detected across all evaluated tasks and models, with F1 scores ranging from 61.1% to 74.5%. This consistency suggests that NeuroFaith captures a systematic aspect of model behavior that manifests similarly across different tasks. Details about class-averaged faithfulness vectors and layer-wise classification score are in Appendix D.3.
Figure 3: Linear vectors max. cosine similarity on gemma-2-27b
Linear Representation Similarity Analysis.
To investigate the hypothesized relationships between faithfulness and related AI safety behaviors, we analyze the representational overlap between NeuroFaith faithfulness vectors and established linear representations of hallucination and deceptiveness from prior work (Rimsky et al., 2024; Goldowsky-Dill et al., 2025). Given that faithful explanations should accurately reflect model reasoning while hallucination and deceptiveness involve misrepresentation of information, we expect these behaviors to exhibit inverse correlations in their neural representations. We restrict analysis to layers with F1 >> 60% where faithfulness is properly linearly represented. Figure 3 shows maximum cosine similarity and corresponding layers for gemma-2-27B (additional results in Appendix D.3). Key findings emerge: Faithfulness vectors are consistently negatively correlated with hallucination and deceptiveness ones (up to -0.38), aligning with expectations about the opposition between faithfulness and both deceptiveness and hallucination. Except Ledgar with 2-hop reasoning, cross-task faithfulness vectors show positive correlation (0.28-0.49), suggesting shared underlying representations of faithful reasoning across domains. These correlations validate that NeuroFaith captures behaviorally meaningful aspects of model reasoning aligned with established AI safety behaviors. Table 6 in Appendix D.3 corroborates these findings, demonstrating cross-task transfer: faithfulness vectors trained on 2-hop reasoning can detect unfaithful self-NLE in AGNews, and AGNews faithfulness vectors can detect unfaithful self-NLE in Ledgar (using gemma-2-27b).
6.2 Faithfulness Enhancement Through Steering on 2-hop reasoning
Having established the linear structure of NeuroFaith faithfulness representations and their negative correlation with hallucination and deceptiveness vectors, we investigate whether linear steering can improve self-NLE faithfulness during inference. We implement activation steering by modifying hidden states during inference: $h^{\ell}_{k}\leftarrow h^{\ell}_{k}+\lambda\times\overrightarrow{SV_{\ell}}$, where $\overrightarrow{SV_{\ell}}$ is the steering vector and $\lambda$ controls intervention intensity. Our objective is here to demonstrate immediate practical value for improving self-NLE faithfulness during inference without model modification. More sophisticated steering methods (Hedström et al., 2025) would likely yield superior results. We evaluate three approaches: faithfulness amplification ($\overrightarrow{SV_{\ell}}=\overrightarrow{F_{\ell}}$, $\lambda=1$) and hallucination/deceptiveness inhibition ($\overrightarrow{SV_{\ell}}=\overrightarrow{H_{\ell}}$ or $\overrightarrow{D_{\ell}}$, $\lambda=-1$), where $\overrightarrow{H_{\ell}}$ and $\overrightarrow{D_{\ell}}$ are hallucination and deceptiveness linear vectors obtained from Rimsky et al. (2024) and Goldowsky-Dill et al. (2025). Faithfulness intervention targets only layers with F1 >> 60%, ensuring modifications occur where faithfulness is reliably encoded.
Figure 4: Faithfulness linear probe visualization examples before ($e$) and after ($e_{steer}$) faithfulness steering. Red: unfaithful activations; Green: faithful activations.
Table 3 shows steering interventions successfully convert 8-11% of unfaithful self-NLE into faithful explanations across all models. Direct faithfulness amplification consistently outperforms hallucination and deceptivenes inhibition, indicating these are distinct behavioral dimensions. Steering effectiveness remains consistent regardless of prediction accuracy, demonstrating genuine explanation quality improvements. Both methods show similar overall gains and highlight the practical utility of NeuroFaith for real-time faithfulness enhancement without model modification. Figure 4 shows how a faithfulness linear probe identifies bridge objects as key tokens determining self-NLE faithfulness, comparing an unfaithful example with its steered faithful version. Additional analyses include comprehensive faithfulness status changes and other examples of converted self-NLE (see Appendix E.1 and H).
Model Faithfulness $\lambda=1$ Hallucination $\lambda=-1$ Deceptiveness $\lambda=-1$
acc. inacc. overall acc. inacc. overall acc. inacc. overall
gemma-2-2b 11.3% 9.3% 11.1% 10.0% 10.1% 10.0% 4.1% 5.7% 4.2%
gemma-2-9b 8.3% 10.8% 8.7% 8.0% 11.2% 8.5% 4.9% 2.5% 4.5%
gemma-2-27b 10.9% 11.3% 11.0% 9.2% 9.1% 9.2% 4.1% 3.2% 3.9%
Table 3: Proportion of initially unfaithful self-NLE made faithful through steering interventions.
7 Conclusion
This work introduces NeuroFaith, a flexible framework that measures self-NLE faithfulness by comparing explanations against mechanistic analysis of model internal activity. NeuroFaith aligns more closely than existing approaches with the common acceptance of faithfulness. Our key contributions suggest that faithfulness exhibits linear structure in LLM representation space, enabling both fast and reliable detection and enhancement through steering interventions. These findings establish concrete pathways toward more transparent AI systems, opening up new avenues of research on faithfulness. NeuroFaith relies on concept extraction methods that may introduce systematic biases, and circuit selection requires domain expertise or expensive circuit discovery that may limit generalizability. Our analysis focuses on predict-then-explain scenarios; extending to explain-then-predict settings could reveal how explanation faithfulness relates to performance gains (Bhan et al., 2024). Additionally, applying NeuroFaith to chain-of-thought reasoning could provide valuable comparisons with existing CoT faithfulness studies (Lanham et al., 2023; Turpin et al., 2023), advancing our understanding of reasoning faithfulness.
Appendix A Appendix
Appendix Table of Contents
A Scientific Libraries
We used several open-source libraries in this work: pytorch (Paszke et al., 2019), HuggingFace transformers (Wolf et al., 2020) and sklearn Pedregosa et al. (2011).
B LLM Implementation Details
Backbone and Special Tokens.
The library used to import the pretrained autoregressive language models is Hugging-Face. In particular, the backbones version of Gemma-2-2B, Gemma-2-9B and Gemma-2-27B are gemma-2-2B-it, gemma-2-9B-it, gemma-2-27B-it respectively. The models were imported with the Bfloat16 computational format. The following special tokens used were used for instruction prompting:
  • user_token= ’<start_of_turn>user’
  • assistant_token= ’<start_of_turn>model’
  • stop_token= ’<eos>’
Text Generation.
Text generation was performed using the native functions of the Hugging Face library: generate. The generate function was used with the following parameters:
  • do_sample = True
  • num_beams = 2
  • no_repeat_ngram_size = 2
  • repetition_penalty =1.2
  • early_stopping = True
  • temperature = 0.05
Self-NLE Generation.
The prompt used to get self-NLE was as follows:
  • <<user>>
  • Question
  • <<Assistant>>
  • Answer
  • <<user>>
  • Give me a simple explanation of your answer.
  • <<Assistant>>
C Datasets
Here we provide detailed information about the analyzed datasets. For 2-hop reasoning, we run NeuroFaith on the Wikidata-2-hop dataset (Biran et al., 2024). We first compute task accuracy on the whole dataset, and then sample 1500 accurate and inaccurate prediction for the following of our study. To sample these 1500 instances, we filter 2-hop reasoning questions where relations are subjective (e.g. ”the most notable work of”) or equivocal, potentially leading to numerous possible answers (e.g. ”the work that features” ). We also sample by setting a maximum number of occurrences for generated answers (15), to foster diversity in the input questions. This filter enables to avoid having to many questions where the answer is a country (e.g. USA). We end up with a dataset made of 3000 samples with 1500 accurate and 1500 inaccurate predictions.
For classification, we run NeuroFaith on AGNews (Gulli, 2005), a newspaper article classification and Ledgar (Tuggener et al., 2020), a more challenging critical domain legal document classification dataset. We retrieve the enriched versions from Bhan et al. (2025) with labeled concepts for CAV computation. For AGNews, the classes to predict are ’world’, ”sport”, ”business” and ”science & technology”. For Ledgar, the classes to predict are ”Amendments”, ”Survival”, ”Terminations” and ”Terms”. Each dataset is made of 4000 samples.
D NeuroFaith Implementation Details
D.1 Concept Extraction
Here we provide the prompts used to give instructions to Qwen-3-32b to extract relevant concepts (NeuroFaith step 1). For 2-hop reasoning, concept extraction consists in retrieving the bridge object from the self-NLE. Since the input text as the following structures : $x=(o_{1},r_{1},{\blacktriangle{}},r_{2},\bullet)$, we directly prompt the model to resolve $(o_{1},r_{1},{\blacktriangle{}}$) by grounding its response on the self-NLE only. We structure our prompt following an in-context learning template with two examples differing from the dataset of interest:
  • <<user>>
  • preprompt + preprompt_example_1
  • <<Assistant>>
  • Emmanuel Macron
  • <<user>>
  • preprompt + preprompt_example_2
  • <<Assistant>>
  • Ingmar Bergman
  • preprompt + $(o_{1},r_{1},{\blacktriangle{}}$)
  • <<user>>
with preprompt = ”Answer briefly and only according to the provided text. If there is no clear answer, say **no bridge object**”, preprompt_example_1 = ”Emmanuel Macron is the president of Italy, and the capital city of Italy is Rome.** the president of Italy is” and preprompt_example_2 = ”The movie Persona is a movie happening in the Faro island and has been directed by Ingmar Bergman, who is from Sweden.**: ’The director of Persona is
For classification, we assess having access to a set of relevant concepts $\mathcal{C}$ and labels related to the task of interest. Given a certain concept $c$, we only prompt the model to assess if the concept is present in the self-NLE if the concept was initially present in the input text, making the concept extraction process computationally less expansive. We structure our prompt following an in-context learning template with three examples differing from the dataset of interest:
  • <<user>>
  • preprompt + preprompt_example_1
  • <<Assistant>>
  • YES
  • <<user>>
  • preprompt + preprompt_example_2
  • <<Assistant>>
  • YES
  • <<user>>
  • preprompt + preprompt_example_3
  • <<Assistant>>
  • NO
  • <<user>>
[…]
with preprompt = ”Analyze whether a given concept has a meaningful impact on predicting a specific category from the provided text explanation. Instructions: (1) Answer with exactly ”YES” if the concept is clearly mentioned in the given text and relevant to the category prediction, (2) Answer with exactly ”NO” if the concept is neither mentioned nor relevant in the given text (3) Consider the logical connection between the concept and the category in the given text”.
For AGNews:
  • preprompt_example_1 = ”Text explanation: ’The article says that OECD countries became richer in the 20th century. This falls under the category of world’. Question: According to the previous text, does the concept ”economic trends” have a meaningful impact on predicting the ”world” category?
  • preprompt_example_2 = ”Text explanation: ’The abstract underlines that the French soccer striker is a good player. It is relevant to sport’. Question: According to the previous text, does the concept ”jargon specific to the sport” have a meaningful impact on predicting the ”sport” category?”.
  • preprompt_example_3 = ”Text explanation: ’The research paper discusses quantum computing algorithms and their complexity. This falls under the category of technology’. Question: According to the previous text, does the concept ”cooking techniques” have a meaningful impact on predicting the ”technology” category?”.
For Ledgar:
  • preprompt_example_1 = ”Text explanation: ’This clause defines what happens to your stock options if your employment ends *before* they are fully vested. It’s part of the core **terms** of your employment agreement that outlines how these shares work.’ Question: According to the previous text, does the legal concept ”Minimum commitment periods” have a meaningful impact on predicting the ”terms” category?
  • preprompt_example_2 = ”Text explanation: ’This clause defines what happens to your stock options if your employment ends *before* they are fully vested. It’s part of the core **terms** of your employment agreement that outlines how these shares work.’ Question: According to the previous text, does the legal concept ”Effect of termination” have a meaningful impact on predicting the ”terms” category?”.
  • preprompt_example_3 = ”Text explanation: ’The clause specifically talks about how changes can be made to the agreement:”terminated, amended, modified or supplemented”. These are all words that mean changing the original terms of the contract. Since it focuses on how the contract itself can be altered, the relevant category is **Amendments**’ Question: According to the previous text, does the legal concept ”Amendment procedures” have a meaningful impact on predicting the ”Amendments” category?”.
[…]
D.2 Mechanistic Interpretation
Here we provide implementation details about mechanistic interpretations of extracted concepts (NeuroFaith step 2).
Circuit Granularity.
To generate an mechanistic interpretation of an LLM hidden state, the first step is to define the specific neural architectural component and granularity $\mathcal{G}$. Given an index $k$ and a layer $\ell$, a granularity $\mathcal{G}$ refers to a specific point within the model’s computational subgraph where neural activations can be analyzed and from which a local neural interpretation $i^{\ell}_{k}(x)$ is going to be generated. Transformer-based generative LLM can be viewed as a stack of decoder computational blocks (Radford et al., 2018). The information flow through a single layer and from a layer to another can be described with the following equations:
$h_{k}^{\ell}$ $=h_{k}^{\ell-1}+a_{k}^{\ell}+m_{k}^{\ell}$
$a_{k}^{\ell}$ $=\text{MHA}^{\ell}\left(h_{1}^{\ell-1},h_{2}^{\ell-1},\ldots,h_{k}^{\ell-1}\right)$
$m_{k}^{\ell}$ $=\text{MLP}^{\ell}(a_{k}^{\ell}+h_{k}^{\ell-1})$ (1)
where $h_{k}^{\ell}$ denotes the residual stream at index $k$ and layer $\ell$, $a_{k}^{\ell}$ the attention output of the multi-head attention operation $\text{MHA}^{\ell}$ and $m_{k}^{\ell}$ the output from the multi-layer perceptron $\text{MLP}^{\ell}$.
This characterization of the information flow through a Transformer block naturally highlights three possible granularities, each offering different perspectives on the model’s information processing: the residual stream (RS) focusing on $h_{k}^{\ell}$, which serves as the main information pathway through the network; the multi-head attention (MHA) focusing on $a_{k}^{\ell}$, which captures token-to-token interactions; and the multi-layer perceptron (MLP) focusing on $m_{k}^{\ell}$ and performing non-linear transformations of the information (Elhage et al., 2021; Rai et al., 2024). These computational blocks offer a relevant granularity to decode the internal activity of an LLM, since focusing on neurons alone can rarely be done due to their polysemanticity (Elhage et al., 2022). Thus, a hidden state granularity is defined such as $\mathcal{G}\in\left\{\texttt{RS, MHA, MLP}\right\}$.
Interpreter.
For the 2-hop reasoning instantiation, the interpreter used to decode $f$ hidden states $\left\{h_{k}^{\ell}\right\}$ is Patchscopes Ghandeharioun et al. (2024). We use the prompt ”What is the following? Answer briefly [X,X]” to generate the interpretation where X is a token placeholder to be replaced in the latent space by the hidden state to be interpreted. We replace the placeholder tokens at layers 3 and 4 to get two interpretations per hidden state to decode. The layers to be interpreted by Patchscopes vary depending on the assessed model. We set the index token $k$ as the last one related to $o_{1}$ and focus on the late early layers as in Biran et al. (2024). This way, $\ell\in\left\{{5,6,7,8,9,10,11}\right\}$ for gemma-2-2B, $\ell\in\left\{{8,9,10,11,12,13,14}\right\}$ for gemma-2-9B and $\ell\in\left\{{11,12,13,14,15,16,17}\right\}$ for gemma-2-27B.
For the classification instantiation and given a concept $c$, concept importance $\texttt{I}_{\Gamma}(c)$ is based on the linear representation of the concept called CAV and denoted $\overrightarrow{c}$. Concepts are selected based on the identifiability of each concept on $f$ representation space based on $\texttt{I}_{\Gamma}(c)$. Below are examples of selected concepts from Ledgar and AGNews for gemma-2-27b with F1 score $60\%$:
Concept Layer F1 Score (%)
Players 36 0.758
Political developments 43 0.842
Scores 10 0.626
Financial markets 35 0.758
Companies 36 0.805
Industry-specific terminology and jargon 34 0.745
Global issues 45 0.812
Sports events 45 0.854
Industry analysis 38 0.610
Economic trends 36 0.798
Industries 36 0.776
International events 36 0.847
Global politics 36 0.809
International relations 36 0.776
Foreign affairs 43 0.795
News about wars, conflicts 43 0.694
Athletic competitions 45 0.855
Teams 45 0.679
Game summaries 34 0.780
Jargon specific to the sport 45 0.854
Charts, graphs, and financial data 38 0.658
Advancements in computing 41 0.607
Technological trends 43 0.769
Table 4: Ledgar concept max. F1 scores ($60\%$) with related layer for gemma-2-27b.
Concept Layer F1 Score (%)
Modification rights 34 0.838
Amendment procedures 39 0.803
Notice requirements 20 0.743
Approval mechanisms 28 0.816
Integration with original agreement 23 0.729
Format requirements 26 0.768
Severability of amendments 37 0.812
Retroactive application 27 0.638
Waiver limitations 30 0.793
Amendment thresholds 42 0.753
Amendment restrictions 36 0.730
Prior versions validity 45 0.759
Amendment documentation 24 0.759
Version control mechanisms 37 0.738
Material change provisions 33 0.754
Post-termination obligations 28 0.798
Duration of surviving terms 28 0.885
Identification of specific clauses 11 0.785
Indemnification continuation 40 0.862
Payment obligations survival 21 0.776
Non-compete/non-solicitation persistence 30 0.631
Representations/warranties survival 21 0.767
Remedies availability post-termination 42 0.845
Perpetual rights 40 0.815
Legal compliance requirements 34 0.721
Duration specifications 42 0.885
Commencement date 36 0.638
Expiration conditions 37 0.910
Renewal mechanisms 44 0.654
Term length 26 0.798
Condition precedents 28 0.798
Milestone-based periods 36 0.739
Initial term vs. renewal term distinctions 44 0.789
Evergreen provisions 13 0.768
Term modification triggers 28 0.728
Minimum commitment periods 34 0.777
Maximum term limitations 36 0.785
Regulatory term constraints 41 0.769
Term acceleration provisions 30 0.738
Rolling term provisions 29 0.823
Termination rights 28 0.888
Notice periods 45 0.726
Termination for convenience 44 0.677
Effect of termination 28 0.912
Wind-down procedures 28 0.865
Early termination penalties 36 0.682
Mutual termination provisions 27 0.854
Partial termination rights 37 0.833
Change of control provisions 29 0.729
Performance-based termination 44 0.719
Regulatory/legal change termination 15 0.776
Termination certification requirements 27 0.774
Post-termination restrictions 13 0.747
Transition obligations 28 0.856
Table 5: Ledgar concept max. F1 scores ($60\%$) with related layer for gemma-2-27b.
[…]
The concept importance $\texttt{I}_{\Gamma}(c)$ can also be computed with gradients-based approaches such as TCAV. It can be formally expressed based on the previously computed CAV $\overrightarrow{c}$. $\texttt{I}_{\Gamma}(c)$ is calculated by aggregating the local importance measures related to the hidden states along circuit $\Gamma$:
$\texttt{I}_{\Gamma}(c)=\underset{(k,\ell)\in\Gamma}{\bigodot}\langle\overrightarrow{c},\nabla f^{\ell}_{\hat{y},k}(h^{\ell}_{k})\rangle$ (2)
[…]
Faithfulness is finally calculated based on concept importance as follows: $F(x,e)=\frac{1}{p}\sum_{i=1}^{p}\mathbf{1}_{\texttt{I}_{\Gamma}(c_{i})>0}$. Below are two distributions of faithfulness for classification of self-NLE generated for AGnews and Ledgar from gemma-2-27b.
In Figure 5, 6, 7, 8, 9, and 10 we plot the concepts sorted by frequency for faithful and unfaithful self-NLE for AGNews and Ledgar for gemma-2-2b and gemma-2-9b for several class predictions. These figures enable to highlight concepts related to either faithful or unfaithful self-NLE. For example, Figure 8 highlights the concept of ”companies” to the class world, which seems to be counterintuitive.
Figure 5: Concepts related to faithful self-NLE and the prediction ”business”, sorted by frequency for AGNews for gemma-2-2b.
Figure 6: Concepts related to unfaithful self-NLE and the prediction ”business”, sorted by frequency for AGNews for gemma-2-2b.
Figure 7: Concepts related to faithful self-NLE and the prediction ”world”, sorted by frequency for AGNews for gemma-2-2b.
Figure 8: Concepts related to unfaithful self-NLE and the prediction ”world”, sorted by frequency for AGNews for gemma-2-2b.
Figure 9: Concepts related to faithful self-NLE and the prediction ”amendments”, sorted by frequency for Ledgar for gemma-2-9b.
Figure 10: Concepts related to unfaithful self-NLE and the prediction ”amendments”, sorted by frequency for Ledgar for gemma-2-9b.
We also plot the density of faithfulness for AGNews and Ledgar for gemma-2-27b in Figure 11 and 12.
Figure 11: NeuroFaith faithfulness distribution for AGNews for gemma-2-27b.
Figure 12: NeuroFaith faithfulness distribution for Ledgar for gemma-2-27b.
D.3 Linear Latent Faithfulness Detection.
Here we provide additional information about layer-wise linear probe performance and similarity between faithfulness vectors and with other AI safety linear vectors. Figures13, 14, 15, 16, 17, 18, 19, 20, 21 show the layer-wise performance of the faithfulness linear probes for 2-hop reasoning, AGNews and Ledgar. As shown in Figure 30, 23 and 31, cosine similarity between task-specific linear faithfulness and AI safety behaviors vectors becomes more pronounced with the size of the model.
Figure 13: Linear faithfulness probe classification performance for 2-hop reasoning, gemma-2-2B.
Figure 14: Linear faithfulness probe classification performance for 2-hop reasoning, gemma-2-9B.
Figure 15: Linear faithfulness probe classification performance for 2-hop reasoning, gemma-2-27B.
Figure 16: Linear faithfulness probe classification performance for AGNews classification, gemma-2-2B.
Figure 17: Linear faithfulness probe classification performance for AGNews classification, gemma-2-9B.
Figure 18: Linear faithfulness probe classification performance for AGNews classification, gemma-2-27B.
Figure 19: Linear faithfulness probe classification performance for Ledgar classification, gemma-2-2B.
Figure 20: Linear faithfulness probe classification performance for Ledgar classification, gemma-2-9B.
Figure 21: Linear faithfulness probe classification performance for Ledgar classification, gemma-2-27B.
(a) Mean operator
(b) Max operator
Figure 22: Linear vectors cosine similarity analysis on gemma-2-2b
(a) Mean operator
(b) Max operator
Figure 23: Linear vectors cosine similarity analysis on gemma-2-9b
(a) Mean operator
(b) Max operator
Figure 24: Linear vectors cosine similarity analysis on gemma-2-27b
Table 6 shows the F1 score for faithfulness linear detection by using the linear representation from one task (e.g. 2-hop reasoning) to predict self-NLE faithfulness from other tasks (e.g. Ledgar and AGNews). These results corroborate the cosine similarity analysis, highlighting that the 2-hop reasoning/AGNews and AGNews/Ledgar pairs are moderately correlated with gemma-2-27b, enabling to properly detect faithfulness (approximately 62%). This phenomenon does not appear on smaller models.
gemma-2b gemma-9b gemma-27b
Base/Target Agnews Ledgar Agnews Ledgar Agnews Ledgar
2-hop reasoning 52.2% 56.1% 47.2% 40.0% 61.3% 36.1%
AGNews 45.3% 42.2% 63.1%
Table 6: Faithfulness detection by transfer, from a base to a target faithfulness linear vector.
E Detailed Taxonomy of Self-NLE in Two-hop Reasoning
Taxonomy Definition.
[…]
Figure 25: Detailed taxonomy of $f$ behavior in two-hop reasoning, based on the status of the prediction, the self-NLE and the latent reasoning.
  • ($C_{1}$) Complete reasoning failure. If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\Gamma,o_{2}\notin\tilde{h^{\ell}_{k}}$. Observation: Wrong prediction, incorrect unfaithful explanation, ground-truth bridge object undetected in circuit $\Gamma$. Interpretation: Evidence suggests failure in first-hop reasoning, with neither explanation nor internal representations containing the expected bridge object.
  • ($C_{2}$) Internal-external reasoning mismatch. If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\exists(k,\ell)\in\Gamma,o_{2}\in\tilde{h^{\ell}_{k}}$. Observation: Wrong prediction, incorrect unfaithful explanation, but ground-truth bridge object detected in circuit $\Gamma$. Interpretation: Model appears to have correct internal knowledge but generates inconsistent explanations, indicating potential reasoning-explanation dissociation because of either deceptiveness or hallucination.
  • ($C_{3}$) Explanation-prediction association. If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$ and $\widehat{o_{2}}=o_{2}$. Observation: Wrong prediction with unfaithful but correct explanation. Interpretation: Model associates correct bridge object with incorrect prediction during explanation generation, suggesting superficial pattern matching without genuinely resolving the first hop of the 2-hop reasoning.
  • ($C_{4}$) First-hop reasoning failure. If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}\neq o_{2}$. Observation: Wrong prediction with faithful but incorrect explanation. Interpretation: Model consistently follows incorrect reasoning pathway, indicating systematic error in first-hop reasoning or concept misunderstanding.
  • ($C_{5}$) Second-hop reasoning failure. If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}=o_{2}$. Observation: Wrong prediction with faithful and correct explanation. Interpretation: Model correctly identifies bridge object but fails in second reasoning step, suggesting error occurs after successful first-hop completion.
  • ($C_{6}$) Shortcut learning. If $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\Gamma,o_{2}\notin\tilde{h^{\ell}_{k}}$. Observation: Correct prediction with unfaithful incorrect explanation, ground-truth bridge object undetected. Interpretation: Evidence suggests direct $o_{1}\rightarrow o_{3}$ association, consistent with shortcut learning behavior that bypasses intermediate reasoning steps.
  • ($C_{7}$) Deceptiveness or hallucination. If $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\exists(k,\ell)\in\Gamma,o_{2}\in\tilde{h^{\ell}_{k}}$. Observation: Correct prediction with unfaithful incorrect explanation, but ground-truth bridge object detected internally. Interpretation: Model possesses correct internal knowledge but generates deceptive (or hallucinated) explanations, suggesting reasoning-explanation dissociation or alternative reasoning pathways. This case is expected to be rare, otherwise highlighting a case where $f$ is not honest in its self-NLE while ”knowing” the ground truth bridge object, raising a problem in $f$ alignment.
  • ($C_{8}$) Explainer parrot. If $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$ and $\widehat{o_{2}}=o_{2}$. Observation: Correct prediction with unfaithful but correct explanation. Interpretation: Model generates expected explanations without corresponding detectable internal reasoning, suggesting post-hoc explanation generation (i.e. ”explainer parrot” behavior).
  • ($C_{9}$) Alternative reasoning pathway. If $\widehat{o_{3}}=o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}\neq o_{2}$. Observation: Correct prediction with faithful but incorrect explanation. Interpretation: Model uses consistent but non-canonical reasoning pathways, indicating systematic bias or alternative reasoning mechanism that leads to correct outcomes through unexpected intermediate steps.
  • ($C_{10}$) Reliable oracle. If $\widehat{o_{3}}=o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}=o_{2}$. Observation: Correct prediction with faithful and correct explanation. Interpretation: Strong evidence for expected canonical reasoning pathway, representing the most interpretable and reliable case for knowledge extraction and model understanding.
While this taxonomy relies on the interpreter ability to decode the model’s internal activity, the systematic patterns observed across different models and tasks provide convergent evidence for these behavioral categories. We give examples of thix taxonomy in Appendix H
Incorrect Predictions
Model C1 C2 C3 C4 C5
Complete Internal-external Explanation- First-hop Second-hop
reasoning failure reasoning mismatch prediction assoc. reasoning failure reasoning failure
gemma-2-2b 23.8% 4.2% 14.4% 16.0% 41.6%
gemma-2-9b 26.1% 4.5% 14.5% 14.1% 40.9%
gemma-2-27b 22.8% 4.4% 12.5% 17.1% 43.1%
Table 7: Distribution of categories for incorrect predictions across model sizes.
Correct Predictions
Model C6 C7 C8 C9 C10
Shortcut Deceptiveness or Explainer Alternative Reliable
learning hallucination parrot reasoning pathway oracle
gemma-2-2b 27.8% 6.1% 17.7% 8.1% 40.3%
gemma-2-9b 14.9% 4.2% 20.0% 6.8% 54.0%
gemma-2-27b 12.8% 3.1% 15.3% 6.4% 62.4%
Table 8: Distribution of categories for correct predictions across model sizes.
E.1 2-hop reasoning detailed results.
Here we detail the results outlined in Section 4 and 6 by breaking down the results at the category-level as introduced above.
Taxonomy Descriptive Analysis.
Table 7 highlights that the categories C1 (complete reasoning failure) and C5 (second-hop reasoning failure) are the most represented across the models. The distributions are overall highly stable for incorrect predictions whereas the category C10 (reliable oracle) increases with the model size for accurate predictions. The category C7 (deceptiveness or hallucination) tends to decrease with model size, but still represents a non negligible part of the self-NLE.
Taxonomy Faithfulness Enhancement Analysis.
Table 9,10 and 11 respectively show the detailed impact of hallucination inhibition, linear faithfulness amplification and deceptiveness inhibition steering on self-NLE faithfulness. Categories C2 and C7 are the most prone to be turned into faithful self-NLE overall. Hallucination and faithfulness steering lead to significantly better results as compared to deceptiveness steering overall. Linear faithfulness amplification gives slightly better results than hallucination inhibition. Hallucination inhibition obtains slightly better results than linear faithfulness amplification for C1 and C6.
New Faithfulness After Steering Through Hallucination Inhibition (%)
Model C1 C2 C3 C6 C7 C8
Complete Internal-external Explanation- Shortcut Deceptiveness or Explainer
reasoning failure reasoning mismatch prediction assoc. learning hallucination parrot
gemma-2-2b 10.8% 33.3% 5.6% 7.4% 36.2% 3.0%
gemma-2-9b 9.2% 13.4% 5.1% 16.1% 35.9% 1.6%
gemma-2-27b 12.0% 22.7% 4.8% 12.5% 48.9% 2.6%
Table 9: Percentage of initially unfaithful explanations that become faithful after hallucination steering interventions, by category and model size.
New Faithfulness After Steering Through Linear Faithfulness Amplification (%)
Model C1 C2 C3 C6 C7 C8
Complete Internal-external Explanation- Shortcut Deceptiveness or Explainer
reasoning failure reasoning mismatch prediction assoc. learning hallucination parrot
gemma-2-2b 8.1% 38.0% 5.1% 9.3% 39.5% 1.1%
gemma-2-9b 6.9% 28.4% 3.7% 10.3% 57.8% 2.0%
gemma-2-27b 9.3% 19.7% 5.3% 10.9% 40.4% 1.3%
Table 10: Percentage of initially unfaithful explanations that become faithful after linear faithfulness steering interventions, by category and model size.
New Faithfulness After Steering Through Deceptiveness Inhibition (%)
Model C1 C2 C3 C6 C7 C8
Complete Internal-external Explanation- Shortcut Deceptiveness or Explainer
reasoning failure reasoning mismatch prediction assoc. learning hallucination parrot
gemma-2-2b 5.1% 3.2% 2.7% 4.7% 13.8% 6.0%
gemma-2-9b 6.1% 11.7% 0.6% 8.0% 18.7% 1.4%
gemma-2-27b 5.0% 6.3% 1.7% 2.7% 9.1% 3.3%
Table 11: Percentage of initially unfaithful explanations that become faithful after deceptiveness steering interventions, by category and model size (third experimental condition).
As shown in Figures 27, 28, and 29, we observe consistent transition patterns when steering makes unfaithful explanations faithful. For initially incorrect predictions, category C1 (complete reasoning failure) consistently transitions to C4 (first-hop reasoning failure), achieving faithfulness approximately 10% of the time under NeuroFaith linear faithfulness amplification and hallucination inhibition. Category C2 (internal-external reasoning mismatch) predominantly transitions to C5 (second-hop reasoning failure) when steering succeeds. Category C3 (explanation-prediction association) exclusively leads to C5 (second-hop reasoning failure) upon becoming faithful.
Initially accurate predictions follow similar patterns. Category C6 (shortcut learning) transitions to C9 (alternative reasoning pathway) when made faithful, while category C8 (explainer parrot) becomes C10 (reliable oracle). Category C7 (deceptiveness or hallucination) can transition to either C9 or C10, though it more commonly becomes a reliable oracle (C10). We give several examples of unfaithful self-NLE made faithful in Appendi H.
(a) Initially inaccurate predictions.
(b) Initially accurate predictions.
Figure 26: Detailed taxonomy transition state analysis, before and after NeuroFaith linear faithfulness steering on gemma-2-2b on 2-hop reasoning.
(a) Initially inaccurate predictions.
(b) Initially accurate predictions.
Figure 27: Detailed taxonomy transition state analysis, before and after hallucination inhibition on gemma-2-2b on 2-hop reasoning.
(a) Initially inaccurate predictions.
(b) Initially accurate predictions.
Figure 28: Detailed taxonomy transition state analysis, before and after NeuroFaith linear faithfulness steering on gemma-2-9b on 2-hop reasoning.
(a) Initially inaccurate predictions.
(b) Initially accurate predictions.
Figure 29: Detailed taxonomy transition state analysis, before and after hallucination inhibition on gemma-2-27b on 2-hop reasoning.
(a) Initially inaccurate predictions.
(b) Initially accurate predictions.
Figure 30: Detailed taxonomy transition state analysis, before and after NeuroFaith linear faithfulness steering on gemma-2-27b on 2-hop reasoning.
(a) Initially inaccurate predictions.
(b) Initially accurate predictions.
Figure 31: Detailed taxonomy transition state analysis, before and after hallucination inhibition on gemma-2-2b on 2-hop reasoning.
F NeuroFaith Faithfulness Measure Comparison
In this section we examine how faithfulness as measured by NeuroFaith compares to existing Counterfactual Intervention (CI) Atanasova et al. (2023) and Attribution Agreement (AA) approaches Parcalabescu & Frank (2024) for evaluating LLM self-NLE faithfulness. We focus on 2-hop reasoning and begin with a qualitative comparison that illustrates potential differences between these approaches, followed by a quantitative analysis that provides evidence for these observations.
F.1 2-hop Reasoning Faithfulness Qualitative Comparison.
To illustrate the critical differences between existing faithfulness evaluation approaches and NeuroFaith, we examine a concrete 2-hop reasoning example (see Figure 32) that illusrrates fundamental limitations in current methods (CI and AA). Given input text $x$, prediction $f(x)$, and self-generated explanation $e(x)$, we compare how different faithfulness approaches evaluate the same case. The Example Setup:
  • Original input: $x$ = ”The father of Carol Chomsky is
  • Original prediction: $f(x)$ = ”Harry Abraham Schatz
  • Counterfactual Intervention: $x_{cf}$ = ”The father of the spouse of Carol Chomsky is
  • Counterfactual prediction: $f(x_{cf})$ = ”William Chomsky
Counterfactual Intervention (CI) would assess the self-NLE $e(x_{cf})$ as faithfulness. This evaluation is based solely on the presence of the intervention term ”spouse” within $e(x_{cf})$, establishing consistency between the input modification and explanation content. Moreover, Attribution Analysis (AA) methods consists in comparing attribution scores (e.g., using SHAP) to highlight important tokens (e.g. ”father”, ”spouse”, ”Carol”, ”Chomsky”) for the prediction and the self-NLE. High correlation coefficients between the attribution vectors for prediction and self-NLE would similarly classify the self-NLE as faithful.
However, NeuroFaith rejects $e(x_{cf})$ as unfaithful because the bridge object (”Morris Halle”) is not contained in the decoded hidden states ($\{\tilde{h^{\ell}_{k}}\}$). This mismatch between internal computation and self-NLE content reveals unfaithfulness. The steered self-NLE obtained with linear faithful steering (see Section 6) is evaluated as faithful by NeuroFaith, due to shared bridge object (”Noam Chomsky”) between the new self-NLE and the decoded hidden states.
This qualitative example illustrates that both CI and AA approaches may employ more lenient evaluation standards compared to NeuroFaith and can miss unfaithful self-NLE. This observation is corroborated by our analysis in the next paragraph.
Figure 32: Qualitative comparison between NeuroFaith and CI faithfulness.
F.2 2-hop Reasoning Faithfulness Quantitative Comparison.
We propose an experimental protocol to evaluate the practical utility of different faithfulness measures for model analysis. The protocol is motivated by two common applications of explanations in AI systems: (1) troubleshooting models by identifying the source of errors in wrong predictions Biecek & Samek (2024), and (2) detecting potential biases or shortcuts in correct predictions to ensure they align with expected reasoning processes Ribeiro et al. (2016).
We test whether faithful explanations, as identified by NeuroFaith and CI methods, provide better diagnostic information for these purposes. The key hypothesis is that if a faithfulness measure accurately captures the model’s internal reasoning, then explanations deemed faithful should better localize reasoning failures and identify non-canonical reasoning pathways.
Experimental Framework.
[…]
  • Category A: Wrong prediction ($\widehat{o}_{3}\neq o_{3}$), wrong bridge object ($\widehat{o}_{2}\neq o_{2}$) $\rightarrow$ likely first-hop failure
  • Category B: Wrong prediction ($\widehat{o}_{3}\neq o_{3}$), correct bridge object ($\widehat{o}_{2}=o_{2}$) $\rightarrow$ likely second-hop failure
  • Category C: Correct prediction ($\widehat{o}_{3}=o_{3}$), wrong bridge object ($\widehat{o}_{2}\neq o_{2}$) $\rightarrow$ alternative reasoning pathway
  • Category D: Correct prediction ($\widehat{o}_{3}=o_{3}$), correct bridge object ($\widehat{o}_{2}=o_{2}$) $\rightarrow$ canonical reasoning
Each category can be further subdivided with respect to faithfulness, enabling comparison between faithful and unfaithful self-NLE within each reasoning pattern. In the following we propose 3 metrics based on these 4 categories and assess if faithful self-NLE lead to better model debugging and bias targeting than unfaithful self-NLE.
Figure 33: Simplified taxonomy of $f$ behavior in two-hop reasoning, based on the status of the prediction and the self-NLE.
First Hop Hint.
We test whether faithful self-NLE better identify first-hop reasoning failures through a targeted intervention. For each input $x$ having led to a wrong prediction ($\widehat{o}_{3}\neq o_{3}$), we create a modified version $x_{hint1}=(o_{1},r_{1},o_{2},x)$ that explicitly provides the correct bridge object. For example:
  • $x=$ ”The country of origin of the movie maker that directed Persona is”
  • $x_{hint1}=$ ”The movie maker that directed Persona is Ingmar Bergman. The country of origin of the movie maker that directed Persona is”
If faithful self-NLE accurately reflect internal reasoning, then providing first-hop hints should differentially improve performance for Category A (first-hop failures) versus Category B (second-hop failures), and this difference should be stronger for faithful self-NLE. We define the performance ratio under first-hop hints as:
$PR(A,B,hint1)=\frac{ACC(A,hint1)}{ACC(B,hint1)}$ (3)
where $ACC(A,hint1)$ represents the accuracy of Category A examples when given first-hop hints. We compute separate ratios for faithful and unfaithful explanations:
$PR(A,B,hint1,faithful)=\frac{ACC(A,hint1,faithful)}{ACC(B,hint1,faithful)}$ (4)
Finally, the Compound Accuracy Score (CAS) quantifies whether faithful explanations provide better error localization:
$CAS(A,B,hint1)=log\Bigl(\frac{PR(A,B,hint1,faithful)}{PR(A,B,hint1,unfaithful)}\Bigr)$ (5)
A positive CAS indicates that faithful self-NLE better identify first-hop failures.
Quality Metric gemma-2-2b gemma-2-9b gemma-2-27b
NeuroFaith CI NeuroFaith CI NeuroFaith CI
hint1 0.04 -0.07 0.06 -1.10 0.75 -0.99
hint2 -0.44 -0.55 0.28 -0.03 -0.35 1.25
$r_{2}\rightarrow r^{\prime}_{2}$ 0.09 -2.62 0.17 -0.32 0.31 -0.26
Table 12: NeuroFaith comparison to CI (higher is better) across models on 352 CI-compatible samples on 2-hop reasoning.
Quality Metric gemma-2-2b gemma-2-9b gemma-2-27b
hint1 0.48 0.14 0.22
hint2 0.03 0.33 0.88
$r_{2}\rightarrow r^{\prime}_{2}$ -0.04 1.18 0.40
Table 13: NeuroFaith evaluation across models on the overall 2-hop reasoning dataset.
Table 14: Correlation analysis between NeuroFaith and CI.
Second Hop Hint.
We now test whether faithful self-NLE better identify second-hop reasoning failures through a targeted intervention. Following the logic introduced above, for each input $x$ having led to a wrong prediction ($\widehat{o}_{3}\neq o_{3}$), we create a modified version $x_{hint2}=(o_{2},r_{2},o_{3},x)$ that explicitly provides the second part of the 2-hop reasoning. For example:
  • $x=$ ”The country of origin of the movie maker that directed Persona is”
  • $x_{hint2}=$ ”The country of origin of Ingmar Bergman is Sweden. The country of origin of the movie maker that directed Persona is”
If faithful self-NLE accurately reflect internal reasoning, then providing second-hop hints should differentially improve performance for Category B (second-hop failures) versus Category A (first-hop failures), and this difference should be stronger for faithful self-NLE. We build our second metric following the notations introduced above, based on the Compound Accuracy Score:
$CAS(B,A,hint2)=log\Bigl(\frac{PR(B,A,hint2,faithful)}{PR(B,A,hint2,unfaithful)}\Bigr)$ (6)
Here, a positive CAS indicates that faithful self-NLE better identify second-hop failures.
Second Relation Modification.
We test whether faithful self-NLE better identify non-canonical reasoning pathways through the modification of the second step of the reasoning trace ($r_{2}$). Intuitively, a non-canonical reasoning pathways is more prone to lead to false reasoning when chaning one step of the 2-hop reasoning. For each input $x$ having led to a good prediction ($\widehat{o}_{3}=o_{3}$), we create a modified version $x_{r_{2}\rightarrow r^{\prime}_{2}}=(o_{1},r_{1},{\blacktriangle{}},r_{2}^{\prime},\bullet)$ that change the second relation of the 2-hop reasoning input. For example:
  • $x=$ ”The country of origin of the movie maker that directed Persona is”
  • $x_{r_{2}\rightarrow r^{\prime}_{2}}=$ ”The father of the movie maker that directed Persona is”
If faithful self-NLE accurately reflect internal reasoning, then changing $r_{2}$ should differentially decrease performance for Category C (alternative reasoning pathway) versus Category D (canonical reasoning), and this difference should be stronger for faithful self-NLE. We build our third metric following the notations introduced above, based on the Compound Accuracy Score:
$CAS(D,C,r_{2}\rightarrow r^{\prime}_{2})=log\Bigl(\frac{PR(D,C,r_{2}\rightarrow r^{\prime}_{2},faithful)}{PR(D,C,r_{2}\rightarrow r^{\prime}_{2},unfaithful)}\Bigr)$ (7)
A positive CAS indicates that faithful self-NLE better identify non-canonical reasoning pathways.
Experimental Results.
We evaluate our three faithfulness indicators using both NeuroFaith and Counterfactual Intervention (CI). The Wikidata-2-hop dataset provides natural support for computing $x_{r_{2}\rightarrow r^{\prime}_{2}}$ instances, as it contains multiple reasoning chains involving similar objects. Additionally, the dataset includes counterfactual interventions through variations in reasoning chains, enabling straightforward CI computation across multiple instances.
Our evaluation reveals three key findings. First, Table 14 shows that NeuroFaith consistently outperforms CI on the subset of CI-compatible samples. Second, Table 14 presents NeuroFaith results on the complete dataset, showing positive metric values across all models except gemma-2-2b.
G Classification Examples
In this section we give two examples of instances characterized as either faithful or unfaithful in classification case.
Figure 34: Example from the AGNews dataset where we detect ”sport event” and ”athletic competition” as relevant concepts from the explanation. These two concepts are assessed as important for the prediction, making this explanation faithful ($F(x,e)=1$).
Figure 35: Example from the Ledgar dataset where we detect ”post-termination obligations”, ”duration of surviving terms” and ”payment obligations survival” as relevant concepts from the explanation. Two concepts over three are assessed as important, giving an explanation faithfulness score at 0.67.
H 2-hop Reasoning Taxonomy Examples
In this section we give examples of characterized instances based on the taxonomy introduced in Appendix E. We also give examples of unfaithful self-NLE made faithful, characterized by the same taxonomy.
Figure 36: Example from the Wikidata-2-hop dataset where we observe an internal-external reasoning mismach. The model incorrectly answers ”Jobete Music” and provides a self-NLE referencing to the bridge object ”Sam Cooke”. This bridge object is false and does not appear in the set of natural language interpretations of $f$ latent states. However, the ground truth bridge object ”Louis Armstrong” is decoded at token index 11 and layers 13 and 14.
Figure 37: Example from the Wikidata-2-hop dataset where we observe shortcut learning. The model correctly answers ”France” and provides a self-NLE referencing to ”Alain Resnais”. This bridge object is incorrect and does not appear in the set of natural language interpretations of $f$ latent states.
Figure 38: Example from the Wikidata-2-hop dataset where we observe shortcut learning. The model correctly answers ”Germany” without providing any bridge object in its self-NLE. The expected bridge object ”The Lost Honour of Katharina Blum” is however decoded from the representation space.
Figure 39: Example from the Wikidata-2-hop dataset where we observe an explainer parrot case. The model correctly answers ”Germany” and provides a self-NLE referencing to ”Wim Wenders”. This bridge object is correct but does not appear in the set of natural language interpretations of $f$ latent states.
Figure 40: Example from the Wikidata-2-hop dataset where we observe an alternative reasoning pathway. The model correctly answers ”Japan” and provides a self-NLE referencing to ”Square Enix”. This bridge object is incorrect and also appears in the set of natural language interpretations of $f$ latent states.
Figure 41: Example from the Wikidata-2-hop dataset where we observe an alternative reasoning pathway. The model correctly answers ”USA” and provides a self-NLE referencing to ”Pierre Omidyar”. This bridge object is correct and also appears in the set of natural language interpretations of $f$ latent states.
Figure 42: Example from the Wikidata-2-hop dataset where we start from an incorrect unfaithful explanation and go to a correct and faithful explanation through NeuroFaith linear faithfulness steering.
Figure 43: Example from the Wikidata-2-hop dataset where we start from an incorrect unfaithful explanation and go to a correct and faithful explanation through NeuroFaith linear faithfulness steering.
Figure 44: Example from the Wikidata-2-hop dataset where we start from an incorrect unfaithful explanation and go to a still incorrect but faithful explanation through NeuroFaith linear faithfulness steering.

### v4

@@
-Did I Faithfully Say What I Thought? Bridging the Gap Between LLMs Neural Activity and Self-Explanations
+NeuroFaith: Evaluating LLM Self-Explanation Faithfulness via Internal Representation Alignment
 Abstract
 Large Language Models (LLMs) can generate plausible free text self-explanations to justify their answers.
-However, these natural language explanations may not accurately reflect the model’s actual reasoning process, indicating a lack of faithfulness.
+However, these natural language explanations may not accurately reflect the model’s actual reasoning process, pinpointing a lack of faithfulness.
 Existing faithfulness evaluation methods rely primarily on behavioral tests or computational block analysis without examining the semantic content of internal neural representations.
@@
 We show the versatility of NeuroFaith across 2-hop reasoning and classification tasks.
-Additionally, a linear faithfulness probe based on NeuroFaith is developed to detect unfaithful self-explanations from representation space and improve faithfulness through steering.
+Additionally, we develop a linear faithfulness probe based on NeuroFaith to detect unfaithful self-explanations from representation space and improve faithfulness through steering.
 NeuroFaith provides a principled approach to evaluating and enhancing the faithfulness of LLM free text self-explanations, addressing critical needs for trustworthy AI systems.
+Machine Learning, ICML
 1 Introduction
-Autoregressive Large Language Models (LLMs) can generate plausible self Natural Language Explanation (self-NLE) to support their answers (Wiegreffe & Marasovic, ; Huang et al., 2023).
-Generating self-NLE consists of prompting the LLM to generate an explanation in a predict-then-explain setting, where the model first generates a response to a question and then produces a self-NLE as a justification.
+Autoregressive Large Language Models (LLMs) can generate plausible self Natural Language Explanation (self-NLE) to support their answers (Wiegreffe and Marasovic, 2021; Huang et al., 2023).
+Generating self-NLE consists of prompting the LLM to output an explanation in a predict-then-explain setting, where the model first generates a response to a question and then produces a self-NLE as a justification.
 Unlike their non-generative predecessors, modern LLMs are trained to generate both answers and free text self-NLE that appear credible despite potentially containing persuasive hallucinations (Sahoo et al., 2024).
 This way, despite their logical and coherent appearance favoring trust in the model (Han et al., 2023), LLM-generated self-NLE turn out to not systematically reflect the actual underlying decision-making process of the model, creating a tension between self-NLE plausibility and faithfulness (Agarwal et al., 2024).
-Faithfulness, as defined by Jacovi & Goldberg (2020), measures ”how accurately the explanation reflects the true reasoning process of the model”, a definition widely adopted in the literature (Lyu et al., 2024) and which we likewise follow throughout this work.
-Unfaithful self-NLE can have serious consequences in critical domains, where explanations that appear plausible but lack faithfulness might lead end-users to over-rely on model predictions and make unfair (Luo et al., 2022) or harmful (Kayser et al., 2024) decisions.
+Figure 1: 2-hop reasoning example with LLM answer and decoded hidden representation ($\tilde{h}$).
+Faithfulness linear probe is applied before ($e$) and after ($e_{steer}$) steering.
+Red: unfaithful activations (e.g.
+Morris Halle); Green: faithful activations (e.g.
+Noam Chomsky).
+Faithfulness, as defined by Jacovi and Goldberg (2020), measures ”how accurately the explanation reflects the true reasoning process of the model”, a definition widely adopted in the literature (Lyu et al., 2024) and which we also follow throughout this work.
+Unfaithful self-NLE can have serious consequences in critical domains (Farah et al., 2023), where explanations that appear plausible but lack faithfulness might lead end-users to over-rely on model predictions and make unfair (Luo et al., 2022) or harmful (Kayser et al., 2024) decisions.
 As the use of LLMs is expanding across diverse fields, the combination of their widespread adoption and the simplicity of generating self-NLE through prompting make evaluating their faithfulness increasingly critical.
-Assessing self-NLE faithfulness presents profound difficulties.
-Unlike more structured explainability methods such as attribution (Wiegreffe et al., 2021) or counterfactual approaches (Madsen et al., 2024), the free-form nature of natural language explanations makes faithfulness evaluation particularly difficult.
+Assessing self-NLE faithfulness presents profound difficulties due to the free-form nature of natural language explanations, unlike more structured explainability methods such as attribution (Wiegreffe et al., 2021) or counterfactual approaches (Madsen et al., 2024).
 Numerous methods have been developed to measure self-explanation faithfulness (see e.g.
@@
 (2024)).
-However, (1) they mostly perform behavioral tests and do not examine LLM internal reasoning processes (Atanasova et al., 2023; Siegel et al., 2024, 2025; Matton et al., 2025), and (2) they identify the computational blocks that contribute to prediction and self-NLE without conducting semantic analysis of the neural representations within these blocks (Wiegreffe et al., 2021; Parcalabescu & Frank, 2024; Yeo et al., 2024).
-These shortcomings have led these methods to be characterized as measuring self-consistency rather than genuine faithfulness (Parcalabescu & Frank, 2024), since they fail to establish direct connections between explanations and the model’s actual reasoning processes.
-To overcome these limitations, we introduce NeuroFaith, a flexible framework for directly measuring LLM self-explanation faithfulness.
+However, (1) they mostly perform behavioral tests and do not examine LLM internal reasoning processes (Atanasova et al., 2023; Siegel et al., 2024, 2025; Matton et al., 2025), and (2) they identify the computational blocks that contribute to prediction and self-NLE without conducting semantic analysis of the neural representations within these blocks (Wiegreffe et al., 2021; Parcalabescu and Frank, 2024; Yeo et al., 2025).
+These shortcomings have led these methods to be characterized as measuring self-consistency rather than genuine faithfulness (Parcalabescu and Frank, 2024), since they fail to establish direct connections between explanations and the model’s reasoning processes.
+To overcome these limitations, we introduce NeuroFaith, a flexible framework for directly measuring LLM self-NLE faithfulness.
 Our main contributions are as follows:
@@
   2. 2.
-    We demonstrate the versatility of NeuroFaith by applying it to both 2-hop reasoning tasks and classification problems.
+    We demonstrate the versatility of NeuroFaith by applying it to both 2-hop reasoning and classification tasks.
   3. 3.
     We develop a linear faithfulness probe based on NeuroFaith to efficiently detect unfaithful LLM self-explanations from representation space and improve faithfulness through steering.
+”Figure 1 illustrates NeuroFaith on a 2-hop reasoning question: despite answering correctly, the model’s initial self-NLE ($e$) is unfaithful.
+Linear steering improves faithfulness by aligning the steered self-NLE ($e_{steer}$) with the decoded internal concept (Noam Chomsky) being required for reasoning solving.
 This paper is organized as follows: Section 2 presents how existing approaches measure the faithfulness of LLM self-NLE.
@@
 Attribution Agreement.
-Another way to measure self-NLE faithfulness is to compute post-hoc Attribution Agreement (AA) between the prediction and the NLE (Parcalabescu & Frank, 2024; Wiegreffe et al., 2021; Yeo et al., 2024).
+Another way to measure self-NLE faithfulness is to compute post-hoc Attribution Agreement (AA) between the prediction and the NLE (Parcalabescu and Frank, 2024; Wiegreffe et al., 2021; Yeo et al., 2025).
 AA methods compute attribution scores for both the model’s predictions and its self-NLE and then measure the correlation between these scores to assess faithfulness.
 Higher correlation values indicate greater faithfulness in the model’s self-NLE.
-AA approaches vary in the post-hoc attribution method employed (gradient-based (Sundararajan et al., 2017), SHAP (Lundberg & Lee, 2017) or activation patching (Meng et al., 2022)).
+AA approaches vary in the post-hoc attribution method employed (gradient-based (Sundararajan et al., 2017), SHAP (Lundberg and Lee, 2017) or activation patching (Meng et al., 2022)).
 While AA methods assess whether the same LLM computational blocks were used when generating both the prediction and the self-NLE, they overlook the semantic content of neural representations, leaving the model’s reasoning process only partially treated.
@@
 In the following, we denote an $L$-layer auto-regressive Transformer-based LLM $f$, a set of input texts X and a text of interest $x\in\textbf{X}$.
-We denote $f(x)$ as the model’s answer and $e(x)$ as the corresponding self-NLE produced by $f$ for input $x$.
+We denote $f(x)$ the model’s answer and $e(x)$ the corresponding self-NLE produced by $f$ for input $x$.
 NeuroFaith is a 3-step framework illustrated in Figure 2, summarized below and detailed in the next subsections.
   1. 1.
-    Concept Extraction. We extract from $e(x)$ a set of concepts $\left\{c_{i}\right\}_{i=1}^{p}$ that influence $f(x)$.
+    Concept Extraction. We extract from $e(x)$ a set of concepts $\left\{c_{i}\right\}_{i=1}^{p}$ quoted as important for $f(x)$.
   2. 2.
@@
     Faithfulness Measurement.
-We compute faithfulness $F(x,e)$ by validating that the concepts $\left\{c_{i}\right\}_{i=1}^{p}$ emphasized in the self-NLE actually have significant mechanistic effects on $f(x)$ as determined by their interpretations $\left\{\texttt{I}_{\Gamma}(c_{i})\right\}_{i=1}^{p}$.
+We compute faithfulness $F(x,e)$ by validating that the detected concepts $\left\{c_{i}\right\}_{i=1}^{p}$ in the self-NLE have mechanistic effects on $f(x)$ as determined by their interpretations $\left\{\texttt{I}_{\Gamma}(c_{i})\right\}_{i=1}^{p}$.
 3.1 Concept Extraction
@@
 Recent research has shown that interpretable binary high-level features, referred to as concepts, appear to be linearly encoded within LLM representation space (Elhage et al., 2022; Park et al., 2024).
-This computational understanding aligns with definitions from cognitive science (Ruiz Luyten & van der Schaar, 2024), where concepts are considered as mental entities essential to thought, enabling information integration and categorization (Goguen, 2005).
+This computational understanding aligns with definitions from cognitive science (Ruiz Luyten and van der Schaar, 2024), where concepts are considered as mental entities essential to thought, enabling information integration and categorization (Goguen, 2005).
 This alignment makes concepts an ideal granularity level for model interpretation.
@@
 Human annotation provides high-quality, domain-expert identification of relevant influential concepts but is costly and may not scale to large datasets.
-LLM-as-a-Judge approaches offer scalability but may introduce systematic biases or miss relevant concepts that human experts would identify.
+LLM-as-a-Judge approaches offer scalability but may miss relevant concepts that human experts would identify.
 We provide detailed examples of prompts used to extract concepts from self-NLE with an auxiliary LLM in Appendix D.1.
-Figure 1: NeuroFaith overview. NeuroFaith (1) extracting concepts from the self-NLE and (2) assessing the mechanistic influence of these concepts to finally (3) measure faithfulness.
+Figure 2: NeuroFaith overview.
+NeuroFaith (1) extracts concepts from the self-NLE (Section 3.1), (2) assesses the mechanistic influence of these concepts (Section 3.2) to finally (3) measure faithfulness (Section 3.3).
 3.2 Concept-wise Mechanistic Interpretation
@@
 This granularity avoids focusing on individual neurons due to their polysemanticity (Elhage et al., 2022).
-For example in Figure 1, the circuit is highlighted in blue, starting at token index 12 and layer 1 and finishes at token index 13 and layer L-1.
-Circuits can be obtained through manual investigation of task-specific patterns (Bereska & Gavves, 2024; Wang et al., 2023; Biran et al., 2024) or automated discovery via activation patching (Meng et al., 2022; Conmy et al., 2023), backward attribution (Ferrando & Voita, 2024), or Transcoders (Dunefsky et al., 2024).
+For example in Figure 2, the circuit is highlighted in blue, starting at token index 12 and layer 1 and finishes at token index 13 and layer L.
+Circuits can be obtained through manual investigation of task-specific patterns (Bereska and Gavves, 2024; Wang et al., 2023; Biran et al., 2024) or automated discovery via activation patching (Meng et al., 2022; Conmy et al., 2023), backward attribution (Ferrando and Voita, 2024), or Transcoders (Dunefsky et al., 2024).
 Additional details about circutis are in Appendix D.2.
 Mechanistic Interpreter.
-To mechanistically interpret a concept $c$ across circuit $\Gamma$, we use an interpreter I to compute an attribution score $\texttt{I}_{\Gamma}(c)$ that evaluates whether $c$ impacts $f(x)$.
+To mechanistically interpret a concept $c$ and its impact on $f(x)$ across circuit $\Gamma$, we use an interpreter I to compute an attribution score $\texttt{I}_{\Gamma}(c)$ that evaluates whether $c$ impacts $f(x)$.
 NeuroFaith implements two approaches, each suited to different objectives.
-Probing methods determine whether $c$ is represented within $f$ hidden states on $\Gamma$.
+Probing methods determine whether $c$ is represented within $f$ hidden states in $\Gamma$.
 This involves either generating natural language interpretations of hidden states (using Selfie (Chen et al., 2024) or Patchscopes (Ghandeharioun et al., 2024)) to detect $c$, or training linear probes (Belinkov, 2022) when $c$ is linearly separable with sufficient labels to train the probe.
@@
 They use gradient-based methods like TCAV (Kim et al., 2018) or Representation Engineering (Zou et al., 2023) to directly manipulate concept-related activations across $\Gamma$ and measure behavioral changes, providing causal evidence for concept importance.
-Section 5 details this procedure for measuring faithfulness in the case of classification.
+Section 5 details this procedure in the case of classification.
 3.3 Faithfulness Measurement
 We consider the self-NLE $e(x)$ as faithful when the extracted concepts $\left\{c_{i}\right\}_{i=1}^{p}$ are demonstrably important for $f$’s internal processing according to $\texttt{I}_{\Gamma}(c)$.
-We propose to define the faithfulness of $e(x)$ as the proportion of concepts it contains that are mechanistically assessed as important, i.e.
-that have a positive attribution score: $F(x,e)=\frac{1}{p}\sum_{i=1}^{p}\mathbf{1}_{\texttt{I}_{\Gamma}(c_{i})>0}$.
+We propose to define the faithfulness of $e(x)$ as the proportion of concepts it contains being mechanistically important, i.e.
+having positive attribution score: $F(x,e)=\frac{1}{p}\sum_{i=1}^{p}\mathbf{1}_{\texttt{I}_{\Gamma}(c_{i})>0}$.
 The faithfulness score has different meanings depending on the chosen interpreter I: Probing-based faithfulness measures the detection rate of explanation concepts within the model’s internal representations along circuit $\Gamma$.
@@
 Importance-based faithfulness estimate causal relevance of explanation concepts for the prediction $f(x)$.
-High importance-based faithfulness indicate that most mentioned concepts actually influence $f$ reasoning process.This framework directly captures concept-level alignment between the self-NLE $e(x)$ and $f$ internal reasoning process.
+High importance-based faithfulness indicate that most mentioned concepts actually influence $f$ reasoning process.
+This framework directly captures concept-level alignment between the self-NLE $e(x)$ and $f$ internal reasoning process.
 NeuroFaith faithfulness scores either indicate that the self-NLE accurately reflects internal processing or that the explanation mentions concepts that are not mechanistically important.
-This way, NeuroFaith directly aligns with the established definition of explanation faithfulness introduced above (Jacovi & Goldberg, 2020), offering a principled evaluation of self-NLE faithfulness.
+This way, NeuroFaith directly aligns with the established definition of explanation faithfulness introduced above (Jacovi and Goldberg, 2020), offering a principled evaluation of self-NLE faithfulness.
+Table 1: 2-hop reasoning accuracy, self-NLE correctness, latent first hop correctness and self-NLE faithfulness across models obtained from NeuroFaith.
+on the the Wikidata-2-hop dataset.
+”(In)accurate” represents (in)correct predictions.
+Model Task Accuracy Self-NLE Latent Hop 1 Self-NLE
+Correctness Correctness Faithfulness
+Accurate Inaccurate Accurate Inaccurate Accurate Inaccurate
+gemma-2-2b 7.1% 58.0% 56.0% 48.3% 47.4% 48.4% 57.6%
+gemma-2-9b 16.6% 74.0% 55.4% 58.5% 44.3% 60.8% 55.0%
+gemma-2-27b 22.3% 77.8% 55.6% 64.9% 46.9% 68.8% 60.2%
+mistral-3-7b 21.2% 70.2% 37.9% 52.9% 29.6% 51.1% 32.6%
 4 The Case of 2-Hop Reasoning
-In the previous section, we defined the high level core principle of NeuroFaith.
-We now instantiate NeuroFaith to 2-hop reasoning using probing-based faithfulness, and precisely characterize self-NLE with respect to both its faithfulness and its correctness.
+In the previous section, we defined the high level core principle of NeuroFaith. We now instantiate NeuroFaith to 2-hop reasoning using probing-based faithfulness.
 4.1 NeuroFaith Instantiation for 2-Hop Reasoning
@@
 The presence of a single bridge object in 2-hop reasoning self-NLE makes probing-based interpretations particularly appropriate.
-For $f$ to correctly answer, it must internally compute the bridge object during the first hop before executing the second hop.
+For $f$ to correctly answer, it must either internally compute the bridge object during the first hop before executing the second hop or leverage shortcuts to directly answer based on $o_{1}$.
 Therefore, detecting the extracted bridge object $\widehat{o_{2}}=c$ in $f$ internal representations is a strong indication to assess whether $c$ is important for the prediction.
 We employ natural language interpretation methods such as Selfie and Patchscopes, rather than linear probes, as they provide higher accuracy and are unsupervised (Ghandeharioun et al., 2024).
-Recent work (Biran et al., 2024; Yang et al., 2024) has shown that the bridge object of 2-hop reasoning was resolved on early layers and can be detected by focusing on the last token corresponding to source object $o_{1}$ and the residual stream (RS).
+Recent work (Biran et al., 2024) has shown that the bridge object of 2-hop reasoning was resolved on early layers and can be detected on the last token representation of source object $o_{1}$ and the residual stream (RS).
 We leverage this information to define circuit $\Gamma$ and generate the natural language description $\tilde{h^{\ell}_{k}}$ of hidden state $h^{\ell}_{k}$.
@@
 Self-NLE and Latent Reasoning Characterization.
-Following the notations introduced in Section 3.3, $e(x)$ is assessed faithful if $c$ is detected in at least one hidden state within $\Gamma$ with the probe I ($\texttt{I}_{\Gamma}(c)$=1).
-Beyond faithfulness, we can use the ground truth object ($o_{2}$) to characterize the correctness of $e(x)$ and the latent first hop respectively: $e(x)$ is said to be correct (self-NLE correctness) if its corresponding bridge object is the expected ground truth ($c=o_{2}$).
+Following the notations introduced in Section 3.3, $e(x)$ is assessed faithful if $c$ is detected in at least one hidden state within $\Gamma$ with the probe I ($\texttt{I}_{\Gamma}(c)$=1), implying that $F(x,e)\in\left\{0,1\right\}$ for 2-hop reasoning.
+Beyond faithfulness (self-NLE faithfulness), we can use the ground truth object ($o_{2}$) to characterize the correctness of $e(x)$ and the latent first hop respectively: $e(x)$ is said to be correct (self-NLE correctness) if its corresponding bridge object is the expected ground truth ($c=o_{2}$).
 Likewise, the model performs correct first-hop latent reasoning (latent hop 1 correctness) if $\exists(k,l)\in\Gamma$ such that $o_{2}\in\tilde{h^{\ell}_{k}}$.
@@
 We give a thorough taxonomy of self-NLE in the case of 2-hop reasoning in Appendix E, based on prediction correctness, self-NLE faithfulness and correctness and latent reasoning correctness.
-We also propose an experimental protocol to evaluate the relevance of our faithfulness measure in Appendix F and compare NeuroFaith to CI, showing our faithfulness measure relevancy in absolute and in comparison to CI.
 4.2 2-Hop Reasoning Experimental Analysis
 Experimental Setup.
-We evaluate NeuroFaith for 2-hop reasoning on the Wikidata-2-hop dataset (Biran et al., 2024).
-We apply NeuroFaith to Gemma-2-2B, Gemma-2-9B and Gemma-2-27B (Riviere et al., 2024).
-We use Qwen3-32B as auxiliary model to extract bridge objects from self-NLE and use the prompt from Ghandeharioun et al.
-(2024) to apply Patchscopes as interpreter.
-We evaluate Latent Reasoning Hop 1 Correctness, Self-NLE Faithfulness and Self-NLE Correctness based on the characterization introduced above.
-The results are computed according to the status of the prediction preceding the self-NLE (accurate vs.
-inaccurate).
-We give more details about the layers used to define $\Gamma$ and the prompt used to extract the bridge object and compute the interpretations in Appendix D.1 and  D.2.
-Model Task Acc. Self-NLE Latent Hop 1 Self-NLE
-Correctness Correctness Faithfulness
-Accurate Inaccurate Accurate Inaccurate Accurate Inaccurate
-gemma-2-2b 7.1% 58.0% 56.0% 48.3% 47.4% 48.4% 57.6%
-gemma-2-9b 16.6% 74.0% 55.4% 58.5% 44.3% 60.8% 55.0%
-gemma-2-27b 22.3% 77.8% 55.6% 64.9% 46.9% 68.8% 60.2%
-Table 1: 2-hop reasoning accuracy, first hop latent reasoning correctness and self-NLE correctness and faithfulness across models obtained from NeuroFaith.
-”(In)accurate” represents the set of predictions initially (in)correct.
+We evaluate NeuroFaith on two 2-hop reasoning datasets: Wikidata-2-hop (Biran et al., 2024) and the more challenging Socrates (Yang et al., 2025).
+For Wikidata-2-hop, we test Gemma-2 (2B, 9B, 27B) (Gemma Team, 2024), and Mistral-7B-Instruct-v0.3 (Jiang et al., 2023); for Socrates, we restrict evaluation to Gemma-2-27B given its increased reasoning difficulty.
+We use Qwen3-32B to extract bridge objects from self-NLEs and Patchscopes (Ghandeharioun et al., 2024) as interpreter.
+We report Latent Hop 1 Correctness, Self-NLE Faithfulness, and Self-NLE Correctness, stratified by prediction accuracy.
+Details on circuit $\Gamma$ and prompts to extract the bridge object and run Patchscopes are provided in Appendices D.1 and D.2.
+Since NeuroFaith relies on Qwen3-32B to extract bridge objects, we evaluate its reliability for this task.
+When prompted to retrieve bridge objects from complete 2-hop reasoning chains ($o_{1}$, $r_{1}$, $o_{2}$) and ($o_{2}$, $r_{2}$, $o_{3}$), the latter achieves 99.5% accuracy.
+We also report results using Phi-4 (Abdin et al., 2024) as an alternative extractor in AppendixD.2, Table 7, showing approximately 90% correlation with Qwen3-32B.
+These findings confirm that NeuroFaith is robust to the choice of auxiliary model and concepts are reliably extracted.
 Key Findings.
@@
 These results support that correct reasoning tend to produce more faithful and correct self-NLE.
-Key failure patterns emerge: approximately 45% of inaccurate predictions correctly resolve the first hop operation, indicating that failure often occurs in the second reasoning step ($o_{2},r_{2},o_{3}$).
+Key failure patterns emerge with Gemma: approximately 45% of inaccurate predictions correctly resolve the first hop operation, indicating that failure often occurs in the second reasoning step ($o_{2},r_{2},o_{3}$).
 This suggests that models can identify correct bridge objects even when failing to complete the full reasoning chain.
-The difference between latent reasoning hop 1 correctness and self-NLE correctness shows that models sometimes identify correct bridge objects internally but fail to express them in their self-NLE.
+The difference between latent hop 1 correctness and self-NLE correctness shows that models sometimes identify correct bridge objects internally but fail to express them in their self-NLE.
 This gap decreases with model size, suggesting better alignment between internal and explicit reasoning in larger models.
+Moreover, self-NLE correctness and faithfulness are comparable between Mistral and Gemma for accurate predictions but substantially lower for Mistral on inaccurate ones.
 Remarkably, even when models produce correct final answers, they identify the correct bridge object in their internal representations only 48-65% of the time.
@@
 The models may be leveraging (1) direct associations between source and target entities (shortcut learning (Geirhos et al., 2020)), or (2) alternative pathways making the model accurately answer for a reason different than the ground truth bridge object (see category 9 in Figure 28).
-This phenomenon highlights that task accuracy alone is insufficient for evaluating reasoning quality.
+Figure 3 shows the results obtained by applying NeuroFaith on Socrates, focusing on predictions with accurate latent first hop reasoning.
+Like the Wikidata-2-hop dataset, accurate predictions lead to more faithful and accurate self-NLE.
+Table 8 in Appendix D.2 shows the complete results obtained from the Socrates dataset.
+We additionally compare NeuroFaith to two commonly used faithfulness measures: CI (Atanasova et al., 2023) and AA (Wiegreffe et al., 2021).
+For this comparison, we use an In-Context Editing evaluation protocol, extending the approach proposed by Zaman and Srivastava (2025).
+Results in Appendix F (Table 16) show that NeuroFaith outperforms these baselines in 6 out of 9 settings.
+Finally, Appendix D.2 presents a sensitivity analysis varying circuit size.
+Figures 13-15 show that faithfulness scores remain consistent across circuit sizes, with only a 10% gap between single-node and full circuits, demonstrating robustness to circuit specification.
+Figure 3: NeuroFaith self-NLE correctness and faithfulness on Socrates on instances with accurate first hop resolution.
 5 The Case of Classification
-This section instantiates NeuroFaith for classification problems by employing concept importance methods as the mechanistic interpreter I.
-For classification tasks, given an input text $x\in\textbf{X}$, we denote $f(x)=\hat{y}\in\mathcal{Y}$ with probability score $p_{\hat{y}}(x)$ where $\mathcal{Y}$ is the label space.
+This section instantiates NeuroFaith for classification by employing concept importance methods as mechanistic interpreter I.
+Given an input text $x\in\textbf{X}$, we denote $f(x)=\hat{y}\in\mathcal{Y}$ with probability score $p_{\hat{y}}(x)$ where $\mathcal{Y}$ is the label space.
 We assume access to a predefined set of task-relevant concepts $\mathcal{C}$ with concept labels for input texts.
@@
 For a concept $c\in\mathcal{C}$, a token index $k$ and a layer $\ell$, the layer-wise CAV is defined as: $\overrightarrow{c_{\ell}}=\frac{1}{\left|\textbf{X}^{+}_{c}\right|}\sum_{x\in\textbf{X}^{+}_{c}}{h^{\ell}_{k}}-\frac{1}{\left|\textbf{X}^{-}_{c}\right|}\sum_{x\in\textbf{X}^{-}_{c}}{h^{\ell}_{k}}$, where $\textbf{X}^{+}_{c}$ and $\textbf{X}^{-}_{c}$ respectively represent the sets of texts from X where the concept $c$ is present or absent and $h^{\ell}_{k}$ denotes the hidden state at layer $\ell$ and token position $k$ at the specified granularity $\mathcal{G}$.
-We set the token index to the final position of $x$, as this location represents the model’s complete computational state prior to next-token generation.
+We set the token index to the final position of $x$, as this location represents $f$’s complete computational state prior to next-token generation.
 These CAVs allow for computing the importance of each concept using Representation Engineering techniques.
@@
 We model this relationship as linear: $p_{\hat{y}}(\cdot)=\beta_{0}+\beta_{1}\times\lambda$, where the concept importance score is $\texttt{I}_{\Gamma}(c)=\beta_{1}$ (i.e.
-the marginal effect of intervention on prediction probability, computed via linear regression across different $\lambda$ values).
+the marginal effect of intervention on probability score).
 $\texttt{I}_{\Gamma}(c)$ is set to 0 when its significance $t$-test shows $p>0.01$.
-This process is repeated for each extracted concept from $e(x)$ to derive faithfulness scores as described in Section 3.3.
-Implementation details and examples are provided in Appendix D.2 and G.
-Model Task Acc. Self-NLE Faithfulness
+This process is repeated for each extracted concept from $e(x)$ to derive faithfulness scores as described in Section 3.3, with $F(x,e)\in\left[0,1\right]$ for classification tasks.
+Details and examples are provided in Appendices D.2 and G.
+Table 2: Classification accuracy and self-NLE faithfulness from NeuroFaith. ”(In)accurate represents (in)correct predictions.
+Model Task Accuracy Self-NLE Faithfulness
 AGNews Ledgar AGNews Ledgar
@@
 gemma-2-27B 88.9% 58.1% 74.0% 72.4% 53.0% 31.6%
-Table 2: Classification accuracy and self-NLE faithfulness across models and datasets obtained from NeuroFaith. ”(In)accurate represents the set of predictions initially (in)correct.
+mistral-3-7B 80.6% 82.9% 86.6% 84.1% 94.8% 84.4%
 5.2 Classification Experimental Analysis
@@
 We evaluate NeuroFaith’s classification instantiation on two datasets with varying complexity: AGNews (Gulli, 2005), a newspaper article classification and Ledgar (Tuggener et al., 2020), a more challenging critical domain legal document classification dataset.
-We apply NeuroFaith to Gemma-2-2B, Gemma-2-9B and Gemma-2-27B (Riviere et al., 2024).
+We apply NeuroFaith to Gemma-2 (2B, 9B, 27B) (Gemma Team, 2024) and Mistral-7B-Instruct-v0.3 (Jiang et al., 2023).
 We use the concept set $\mathcal{C}$ and labels of AGNews and Ledgar from Bhan et al.
@@
 Table 2 shows experimental results for AGNews and Ledgar datasets.
-Self-NLE faithfulness is consistently higher on the simpler AGNews dataset, aligning with the intuition that explanation faithfulness might decrease with task complexity.
+Higher average self-NLE faithfulness is observed with Gemma on AGNews and Mistral on Ledgar, supporting that explanation faithfulness might correlate with task accuracy.
 Accurate predictions generally yield more faithful explanations, suggesting that correct predictions might rely on relevant concepts.
-Notably, gemma-2-9B achieves the highest faithfulness scores across both datasets, indicating better alignment and going against the idea of a monotonic relationship between model scale and explanation faithfulness.
-Comparing findings across classification and 2-hop reasoning reveals both general patterns and task-specific differences.
-General findings: (1) accurate predictions tend to produce more faithful explanations, and (2) increased task difficulty seems to reduce self-NLE faithfulness.
-However, scaling effects vary between tasks, indicating that the model scale-faithfulness relationship is highly task-dependent, a finding consistent with prior work (Madsen et al., 2024; Atanasova et al., 2023; Parcalabescu & Frank, 2024; Matton et al., 2025).
+Notably, gemma-2-9B and Mistral-3-7B respectively achieve the highest faithfulness scores on AGNews and Ledgar, going against the idea of a monotonic relationship between model scale and explanation faithfulness.
+Figures 8 and 10 in Appendix E.1 present examples of concepts frequently associated with unfaithful self-NLEs, revealing confusion where the model invokes concepts related to ”business” when predicting the ”world” class, and vice versa.
+Comparing findings across classification and 2-hop reasoning reveals that accurate predictions tend to produce more faithful explanations, whereas scaling effects vary between tasks and models, as shown in prior work (Atanasova et al., 2023; Parcalabescu and Frank, 2024; Matton et al., 2025).
 6 Linearly Detecting and Improving Self-NLE Faithfulness
-Recent work has demonstrated that various AI safety behaviors naturally associated with faithfulness are encoded as linear directions in LLM representation spaces (Bereska & Gavves, 2024), such as hallucination (Rimsky et al., 2024) or deceptiveness (Goldowsky-Dill et al., 2025).
-We investigate whether NeuroFaith-based faithfulness exhibits linear structure enabling (1) accurate detection in representation space and (2) manipulation to improve faithfulness through activation steering.
+Recent work has demonstrated that various safety behaviors naturally associated with faithfulness are encoded as linear directions in LLM representation spaces (Bereska and Gavves, 2024), such as hallucination (Rimsky et al., 2024) or deceptiveness (Goldowsky-Dill et al., 2025).
+We demonstrate that NeuroFaithfaithfulness also exhibits linear structure for (1) accurate detection in representation space and (2) manipulation to improve faithfulness through activation steering.
 6.1 Linear Detection of Faithfulness
-Figure 2: Majority vote faithfulness linear probe performance across models and datasets.
+Figure 4: Faithfulness linear probe performance per model/ dataset.
+Table 3: Proportion of initially unfaithful self-NLE made faithful through steering interventions.
+Model Faithfulness $\lambda=1$ Hallucination $\lambda=-1$ Deceptiveness $\lambda=-1$
+accurate inaccurate overall accurate inaccurate overall accurate inaccurate overall
+gemma-2-2b 11.3% 9.3% 11.1% 10.0% 10.1% 10.0% 4.1% 5.7% 4.2%
+gemma-2-9b 8.3% 10.8% 8.7% 8.0% 11.2% 8.5% 4.9% 2.5% 4.5%
+gemma-2-27b 10.9% 11.3% 11.0% 9.2% 9.1% 9.2% 4.1% 3.2% 3.9%
 Faithfulness Linear Representation Computation.
@@
 Unlike 2-hop reasoning tasks, classification self-NLE faithfulness yields continuous scores between 0 and 1.
-To ensure clear class separation, we focus on polarized cases where $F(x)\in\left\{0,1\right\}$ for clear separation.
-We construct input sequences as $x_{nle}=[x,\hat{y},e(x)]$ and extract hidden states from final token positions across all layers.
+To ensure clear class separation, we focus on polarized cases where $F(x,e)\in\left\{0,1\right\}$ for clear separation.
+We construct input sequences as $x_{nle}=[x,f(x),e(x)]$ and extract hidden states from final token positions across all layers.
 We train diff-mean linear probes on these states, yielding layer-wise faithfulness representations $\overrightarrow{F_{\ell}}$.
@@
 Details about class-averaged faithfulness vectors and layer-wise classification score are in Appendix D.3.
-Figure 3: Linear vectors max. cosine similarity on gemma-2-27b
+Table 4: Linear vectors max. cosine similarity on gemma-2-27b with related layer. * indicates statistical significance, with $p<1\%$
+AGNews Ledgar 2 hop Halluc.
+[…]
+[…]
+[…]
+[…]
 Linear Representation Similarity Analysis.
-To investigate the hypothesized relationships between faithfulness and related AI safety behaviors, we analyze the representational overlap between NeuroFaith faithfulness vectors and established linear representations of hallucination and deceptiveness from prior work (Rimsky et al., 2024; Goldowsky-Dill et al., 2025).
-Given that faithful explanations should accurately reflect model reasoning while hallucination and deceptiveness involve misrepresentation of information, we expect these behaviors to exhibit inverse correlations in their neural representations.
-We restrict analysis to layers with F1 >> 60% where faithfulness is properly linearly represented.
-Figure 3 shows maximum cosine similarity and corresponding layers for gemma-2-27B (additional results in Appendix D.3).
-Key findings emerge: Faithfulness vectors are consistently negatively correlated with hallucination and deceptiveness ones (up to -0.38), aligning with expectations about the opposition between faithfulness and both deceptiveness and hallucination.
-Except Ledgar with 2-hop reasoning, cross-task faithfulness vectors show positive correlation (0.28-0.49), suggesting shared underlying representations of faithful reasoning across domains.
-These correlations validate that NeuroFaith captures behaviorally meaningful aspects of model reasoning aligned with established AI safety behaviors.
-Table 6 in Appendix D.3 corroborates these findings, demonstrating cross-task transfer: faithfulness vectors trained on 2-hop reasoning can detect unfaithful self-NLE in AGNews, and AGNews faithfulness vectors can detect unfaithful self-NLE in Ledgar (using gemma-2-27b).
+To investigate the hypothesized relationships between faithfulness and related safety behaviors, we analyze the similarity between NeuroFaith faithfulness vectors and established linear representations of hallucination and deceptiveness from  Rimsky et al.
+(2024) and Goldowsky-Dill et al.
+(2025).
+Given that faithful self-NLE should reflect model reasoning while hallucination and deceptiveness involve misrepresentation of information, we expect their neural representations to exhibit inverse correlations.
+We focus on layers with F1 >> 60% where faithfulness is properly linearly represented.
+Table 4 shows maximum cosine similarity and corresponding layers for gemma-2-27B (additional results in Appendix D.3).
+Faithfulness vectors are in general negatively correlated with hallucination and deceptiveness ones (up to -0.38), aligning with expectations about the opposition between faithfulness and both deceptiveness and hallucination.
+Cross-task faithfulness vectors are overall positively correlated (up to 0.49), suggesting shared underlying representations of faithful across tasks.
+These correlations validate that NeuroFaith captures LLMs behavioral aspects aligned with established safety behaviors.
+These results are corroborated by Table 9 in Appendix D.3, showing cross-task transfer with gemma-2-27b: faithfulness vectors from 2-hop reasoning can detect unfaithful self-NLE in AGNews, and AGNews faithfulness vectors can detect unfaithful self-NLE in Ledgar.
 6.2 Faithfulness Enhancement Through Steering on 2-hop reasoning
@@
 Our objective is here to demonstrate immediate practical value for improving self-NLE faithfulness during inference without model modification.
-More sophisticated steering methods (Hedström et al., 2025) would likely yield superior results.
-We evaluate three approaches: faithfulness amplification ($\overrightarrow{SV_{\ell}}=\overrightarrow{F_{\ell}}$, $\lambda=1$) and hallucination/deceptiveness inhibition ($\overrightarrow{SV_{\ell}}=\overrightarrow{H_{\ell}}$ or $\overrightarrow{D_{\ell}}$, $\lambda=-1$), where $\overrightarrow{H_{\ell}}$ and $\overrightarrow{D_{\ell}}$ are hallucination and deceptiveness linear vectors obtained from Rimsky et al.
-(2024) and Goldowsky-Dill et al.
-(2025).
+More sophisticated steering methods (Hedström et al., 2025; Vogels et al., 2025) would likely yield superior results.
+We evaluate three approaches: faithfulness amplification ($\overrightarrow{SV_{\ell}}=\overrightarrow{F_{\ell}}$, $\lambda=1$) and hallucination/deceptiveness inhibition ($\overrightarrow{SV_{\ell}}=\overrightarrow{H_{\ell}}$ or $\overrightarrow{D_{\ell}}$, $\lambda=-1$), where $\overrightarrow{H_{\ell}}$ and $\overrightarrow{D_{\ell}}$ are hallucination and deceptiveness linear vectors obtained from (Rimsky et al., 2024) and (Goldowsky-Dill et al., 2025).
 Faithfulness intervention targets only layers with F1 >> 60%, ensuring modifications occur where faithfulness is reliably encoded.
-Figure 4: Faithfulness linear probe visualization examples before ($e$) and after ($e_{steer}$) faithfulness steering. Red: unfaithful activations; Green: faithful activations.
 Table 3 shows steering interventions successfully convert 8-11% of unfaithful self-NLE into faithful explanations across all models.
-Direct faithfulness amplification consistently outperforms hallucination and deceptivenes inhibition, indicating these are distinct behavioral dimensions.
-Steering effectiveness remains consistent regardless of prediction accuracy, demonstrating genuine explanation quality improvements.
-Both methods show similar overall gains and highlight the practical utility of NeuroFaith for real-time faithfulness enhancement without model modification.
+Direct faithfulness amplification slightly outperforms hallucination inhibition and substantially outperforms deceptiveness inhibition.
+Steering impact on faithfulness remains consistent regardless of prediction accuracy, demonstrating practical utility of NeuroFaith for real-time faithfulness enhancement without model modification.
 Figure 1 shows how a faithfulness linear probe identifies bridge objects as key tokens determining self-NLE faithfulness, comparing an unfaithful example with its steered faithful version.
 Additional analyses include comprehensive faithfulness status changes and other examples of converted self-NLE (see Appendix E.1 and H).
-Model Faithfulness $\lambda=1$ Hallucination $\lambda=-1$ Deceptiveness $\lambda=-1$
-acc. inacc. overall acc. inacc. overall acc. inacc. overall
-gemma-2-2b 11.3% 9.3% 11.1% 10.0% 10.1% 10.0% 4.1% 5.7% 4.2%
-gemma-2-9b 8.3% 10.8% 8.7% 8.0% 11.2% 8.5% 4.9% 2.5% 4.5%
-gemma-2-27b 10.9% 11.3% 11.0% 9.2% 9.1% 9.2% 4.1% 3.2% 3.9%
-Table 3: Proportion of initially unfaithful self-NLE made faithful through steering interventions.
 7 Conclusion
-This work introduces NeuroFaith, a flexible framework that measures self-NLE faithfulness by comparing explanations against mechanistic analysis of model internal activity.
+This work introduces NeuroFaith, a framework that measures self-NLE faithfulness by measuring alignment between explanations and mechanistic analysis of model internals.
 NeuroFaith aligns more closely than existing approaches with the common acceptance of faithfulness.
-Our key contributions suggest that faithfulness exhibits linear structure in LLM representation space, enabling both fast and reliable detection and enhancement through steering interventions.
-These findings establish concrete pathways toward more transparent AI systems, opening up new avenues of research on faithfulness.
-NeuroFaith relies on concept extraction methods that may introduce systematic biases, and circuit selection requires domain expertise or expensive circuit discovery that may limit generalizability.
-Our analysis focuses on predict-then-explain scenarios; extending to explain-then-predict settings could reveal how explanation faithfulness relates to performance gains (Bhan et al., 2024).
-Additionally, applying NeuroFaith to chain-of-thought reasoning could provide valuable comparisons with existing CoT faithfulness studies (Lanham et al., 2023; Turpin et al., 2023), advancing our understanding of reasoning faithfulness.
-Appendix A Appendix
+Our findings suggest that faithfulness exhibits linear structure in LLM representation space for 2-hop reasoning and classification tasks, with emergent similarity with other established safety behaviors.
+Faithfulness linear structure enables both fast detection and enhancement through steering interventions, opening new avenues of research on faithfulness.
+In this paper, NeuroFaith has been applied on circuit either previously defined or along the last token, based on probe accuracy.
+It would be valuable to run NeuroFaith based on circuit discovery methods.
+Our analysis focuses on predict-then-explain scenarios, extending to explain-then-predict settings could reveal how explanation faithfulness relates to performance gains (Bhan et al., 2024).
+Extending NeuroFaith to chain-of-thought reasoning could also provide valuable comparisons with existing CoT faithfulness studies (Lanham et al., 2023; Turpin et al., 2023).
+Impact Statements
+This work aims to advance the trustworthiness of AI systems by developing methods to evaluate whether LLM self-explanations actually reflect their internal reasoning processes.
+Faithful self-explanations are critical for deploying LLMs in high-stakes domains such as healthcare or legal decision-making, where users must understand why a model reached a specific outcome.
+NeuroFaith provides a principled tool for auditing explanation quality, potentially reducing over-reliance on plausible but misleading justifications.
+However any potential NeuroFaith user must consider faithfulness score with caution, avoiding over-reliance on a single metric to take high-stake decisions.
+We especially higlight the importance of human validation during critical NeuroFaith steps such as concept extraction.
+Appendix
 Appendix Table of Contents
+  • A.
+    Scientific Libraries .....................................................................................................................................................................................A
+  • B.
+    LLM Implementation Details .....................................................................................................................................................................................B
+  • C.
+    Datasets .....................................................................................................................................................................................C
+  • D.
+    NeuroFaith Implementation Details .....................................................................................................................................................................................D
+    • D.1
+      Concept Extraction .....................................................................................................................................................................................D.1
+    • D.2
+      Mechanistic Interpretation .....................................................................................................................................................................................D.2
+    • D.3
+      Linear Latent Faithfulness Detection .....................................................................................................................................................................................D.3
+  • E.
+    Detailed Taxonomy of Self-NLE in Two-hop Reasoning .....................................................................................................................................................................................E
+    • E.1
+      2-hop reasoning detailed results.
+.....................................................................................................................................................................................E.1
+  • F.
+    NeuroFaith Faithfulness Measure Comparison .....................................................................................................................................................................................F
+    • F.1
+      2-hop Reasoning Faithfulness Qualitative Comparison .....................................................................................................................................................................................F.2
+    • F.2
+      2-hop Reasoning Faithfulness Quantitative Comparison .....................................................................................................................................................................................F.2
+  • G.
+    Classification Examples .....................................................................................................................................................................................G
+  • H.
+    2-hop Reasoning Taxonomy Examples .....................................................................................................................................................................................H
 A Scientific Libraries
-We used several open-source libraries in this work: pytorch (Paszke et al., 2019), HuggingFace transformers (Wolf et al., 2020) and sklearn Pedregosa et al. (2011).
+We used several open-source libraries in this work: pytorch (Paszke et al., 2019), HuggingFace transformers (Wolf et al., 2020) and sklearn (Pedregosa et al., 2011).
 B LLM Implementation Details
@@
 Here we provide detailed information about the analyzed datasets.
-For 2-hop reasoning, we run NeuroFaith on the Wikidata-2-hop dataset (Biran et al., 2024).
-We first compute task accuracy on the whole dataset, and then sample 1500 accurate and inaccurate prediction for the following of our study.
-To sample these 1500 instances, we filter 2-hop reasoning questions where relations are subjective (e.g.
+For 2-hop reasoning, we run NeuroFaith on the Wikidata-2-hop dataset (Biran et al., 2024) and Socrates (Yang et al., 2025).
+We first compute task accuracy on the whole dataset, and then sample 1500 accurate and inaccurate prediction for Wikidata-2-hop.
+To sample these 1500 instances, we filter out 2-hop reasoning questions where relations are subjective (e.g.
 ”the most notable work of”) or equivocal, potentially leading to numerous possible answers (e.g.
@@
 We also sample by setting a maximum number of occurrences for generated answers (15), to foster diversity in the input questions.
-This filter enables to avoid having to many questions where the answer is a country (e.g.
+This filter enables to avoid having too many questions where the answer is a country (e.g.
 USA).
@@
 For classification, we run NeuroFaith on AGNews (Gulli, 2005), a newspaper article classification and Ledgar (Tuggener et al., 2020), a more challenging critical domain legal document classification dataset.
-We retrieve the enriched versions from Bhan et al.
-(2025) with labeled concepts for CAV computation.
+We retrieve the enriched versions from (Bhan et al., 2025) with labeled concepts for CAV computation.
 For AGNews, the classes to predict are ’world’, ”sport”, ”business” and ”science & technology”.
@@
 For 2-hop reasoning, concept extraction consists in retrieving the bridge object from the self-NLE.
-Since the input text as the following structures : $x=(o_{1},r_{1},{\blacktriangle{}},r_{2},\bullet)$, we directly prompt the model to resolve $(o_{1},r_{1},{\blacktriangle{}}$) by grounding its response on the self-NLE only.
+Since the input text as the following structures : $x=(o_{1},r_{1},\tiny{\blacktriangle{}},r_{2},\bullet)$, we directly prompt the model to resolve $(o_{1},r_{1},\tiny{\blacktriangle{}}$) by grounding its response on the self-NLE only.
 We structure our prompt following an in-context learning template with two examples differing from the dataset of interest:
-  • <<user>>
+Concept Extraction Prompt for 2-hop Reasoning user
 preprompt + preprompt_example_1
-  • <<Assistant>>
-  • Emmanuel Macron
-  • <<user>>
-  • preprompt + preprompt_example_2
-  • <<Assistant>>
-  • Ingmar Bergman
-  • preprompt + $(o_{1},r_{1},{\blacktriangle{}}$)
-  • <<user>>
+assistant Emmanuel Macron user preprompt + preprompt_example_2 assistant Ingmar Bergman user preprompt + $(o_{1},r_{1},\tiny{\blacktriangle{}}$)
 with preprompt = ”Answer briefly and only according to the provided text.
@@
 For classification, we assess having access to a set of relevant concepts $\mathcal{C}$ and labels related to the task of interest.
-Given a certain concept $c$, we only prompt the model to assess if the concept is present in the self-NLE if the concept was initially present in the input text, making the concept extraction process computationally less expansive.
+Given a certain concept $c$, we only prompt the model to assess if the concept is present in the self-NLE if the concept was initially present in the input text, making the concept extraction process computationally less expensive.
 We structure our prompt following an in-context learning template with three examples differing from the dataset of interest:
-  • <<user>>
+Concept Extraction Prompt for Classification user
 preprompt + preprompt_example_1
-  • <<Assistant>>
-  • YES
-  • <<user>>
-  • preprompt + preprompt_example_2
-  • <<Assistant>>
-  • YES
-  • <<user>>
-  • preprompt + preprompt_example_3
-  • <<Assistant>>
-  • NO
-  • <<user>>
 […]
@@
 Given an index $k$ and a layer $\ell$, a granularity $\mathcal{G}$ refers to a specific point within the model’s computational subgraph where neural activations can be analyzed and from which a local neural interpretation $i^{\ell}_{k}(x)$ is going to be generated.
-Transformer-based generative LLM can be viewed as a stack of decoder computational blocks (Radford et al., 2018).
+Transformer-based generative LLM can be viewed as a stack of decoder computational blocks (Elhage et al., 2022).
 The information flow through a single layer and from a layer to another can be described with the following equations:
@@
 Interpreter.
-For the 2-hop reasoning instantiation, the interpreter used to decode $f$ hidden states $\left\{h_{k}^{\ell}\right\}$ is Patchscopes Ghandeharioun et al.
-(2024).
+For the 2-hop reasoning instantiation, the interpreter used to decode $f$ hidden states $\left\{h_{k}^{\ell}\right\}$ is Patchscopes (Ghandeharioun et al., 2024).
 We use the prompt ”What is the following?
@@
 The layers to be interpreted by Patchscopes vary depending on the assessed model.
-We set the index token $k$ as the last one related to $o_{1}$ and focus on the late early layers as in Biran et al.
-(2024).
+We set the index token $k$ as the last one related to $o_{1}$ and focus on the late early layers as in (Biran et al., 2024).
 This way, $\ell\in\left\{{5,6,7,8,9,10,11}\right\}$ for gemma-2-2B, $\ell\in\left\{{8,9,10,11,12,13,14}\right\}$ for gemma-2-9B and $\ell\in\left\{{11,12,13,14,15,16,17}\right\}$ for gemma-2-27B.
@@
 Technological trends 43 0.769
-Table 4: Ledgar concept max. F1 scores ($60\%$) with related layer for gemma-2-27b.
+Table 5: AGNews concept max. F1 scores ($60\%$) with related layer for gemma-2-27b.
 Concept Layer F1 Score (%)
@@
 Payment obligations survival 21 0.776
-Non-compete/non-solicitation persistence 30 0.631
 Representations/warranties survival 21 0.767
@@
 Wind-down procedures 28 0.865
-Early termination penalties 36 0.682
 Mutual termination provisions 27 0.854
@@
 Table 6: Ledgar concept max. F1 scores ($60\%$) with related layer for gemma-2-27b.
-[…]
+The concept importance $\texttt{I}_{\Gamma}(c)$ can be either computed through representation engineering and concept erasure or gradient-based approaches such as TCAV.
+We perform concept erasure for $\lambda\in[0,1]$ with a step size of 0.1.
+Table 5 and 6 show the F1 scores of the properly detected concepts in gemma-2-27b representation space.
+We estimate the attribution of $c$ given circuit $\Gamma$ as the $\beta_{1}$ parameter of the following regression: $p_{\hat{y}}(x,\left\{\text{do}(H^{\ell}_{k}=h^{\ell}_{k}-\lambda\times\overrightarrow{c_{\ell}})\right\}_{(k,\ell)\in\Gamma})=\beta_{0}+\beta_{1}\times\lambda$.
+The t-test for $\beta_{1}$ significancy is expressed as follows: $t=\frac{\hat{\beta}_{1}}{SE(\hat{\beta}_{1})}=\frac{\hat{\beta}_{1}}{\sqrt{\frac{MSE}{\sum_{i=1}^{n}(\lambda_{i}-\bar{\lambda})^{2}}}}$, where $\lambda_{i}$ is the $i$ realization of $\lambda$ and $\bar{\lambda}$ is the average $\lambda$ on the analyzed sample.
 The concept importance $\texttt{I}_{\Gamma}(c)$ can also be computed with gradients-based approaches such as TCAV.
@@
 Faithfulness is finally calculated based on concept importance as follows: $F(x,e)=\frac{1}{p}\sum_{i=1}^{p}\mathbf{1}_{\texttt{I}_{\Gamma}(c_{i})>0}$.
-Below are two distributions of faithfulness for classification of self-NLE generated for AGnews and Ledgar from gemma-2-27b.
-In Figure 5, 6, 7, 8, 9, and 10 we plot the concepts sorted by frequency for faithful and unfaithful self-NLE for AGNews and Ledgar for gemma-2-2b and gemma-2-9b for several class predictions.
-These figures enable to highlight concepts related to either faithful or unfaithful self-NLE.
-For example, Figure 8 highlights the concept of ”companies” to the class world, which seems to be counterintuitive.
+Figures 5 and 6 show the faithfulness distributions of NeuroFaith for AGNews and Ledgar for gemma-2-27b.
+Figure 5: NeuroFaith faithfulness distribution for AGNews for gemma-2-27b.
+Figure 6: NeuroFaith faithfulness distribution for Ledgar for gemma-2-27b.
+In Figure 7-12 we plot the concepts sorted by frequency for faithful and unfaithful self-NLE for AGNews and Ledgar for gemma-2-2b and gemma-2-9b for several class predictions.
+These figures reveal which concepts are associated with faithful versus unfaithful self-NLEs.
+For instance, Figure 10 shows that NeuroFaith correctly identifies counterintuitive associations (e.g., the concept ”companies” linked to the class ”world”) as unfaithful, demonstrating its reliability.
 Figure 7: Concepts related to faithful self-NLE and the prediction ”business”, sorted by frequency for AGNews for gemma-2-2b.
@@
 Figure 12: Concepts related to unfaithful self-NLE and the prediction ”amendments”, sorted by frequency for Ledgar for gemma-2-9b.
-We also plot the density of faithfulness for AGNews and Ledgar for gemma-2-27b in Figure 11 and 12.
-Figure 11: NeuroFaith faithfulness distribution for AGNews for gemma-2-27b.
-Figure 12: NeuroFaith faithfulness distribution for Ledgar for gemma-2-27b.
+We also plot the density of faithfulness for AGNews and Ledgar for gemma-2-27b in Figure 5 and 6.
+2-hop Reasoning Additional Results.
+We show in Table 7 the results obtained by applying NeuroFaith in the case of 2-hop reasoning with Phi-4 as bridge object extractor.
+The results are highly similar to the ones obtained by using Qwen-3-32B as bridge object extractor (see column ”Faithfulness Corr w/ qwen”).
+Model Judge Self-NLE Latent Hop 1 Self-NLE Faithfulness
+Correctness Correctness Faithfulness Corr w/ qwen
+Accurate Inaccurate Accurate Inaccurate Accurate Inaccurate
+gemma-2-2b Phi-4 57.80% 56.50% 47.50% 48.30% 48.00% 56.50% 86.50%
+gemma-2-9b Phi-4 73.70% 55.10% 58.50% 44.30% 61.70% 54.80% 91.20%
+gemma-2-27b Phi-4 77.80% 55.70% 64.90% 46.90% 69.10% 59.80% 90.90%
+Table 7: Self-NLE correctness, first hop latent reasoning correctness, and self-NLE faithfulness across models, evaluated using Phi-4 as judge.
+”(In)accurate” represents the set of predictions initially (in)correct.
+”Faithfulness Corr w/ qwen” shows faithfulness correlation between faithfulness scores obtained based on Phi-4 and Qwen-3-32B.
+Model Task Acc. Self-NLE Latent Hop 1 Self-NLE
+Correctness Correctness Faithfulness
+Accurate Inaccurate Accurate Inaccurate Accurate Inaccurate
+gemma-2-27B 3.4% 72.8% 62.1% 36.4% 8.1% 27.3% 7.6%
+Table 8: Self-NLE correctness, first hop latent reasoning correctness, and self-NLE faithfulness for gemma-2-27B on the Socrates dataset.
+Figure 13: NeuroFaith faithfulness sensitivity analysis with regards to circuit size, gemma-2-2B.
+Figure 14: NeuroFaith faithfulness sensitivity analysis with regards to circuit size, gemma-2-9B.
+Figure 15: NeuroFaith faithfulness sensitivity analysis with regards to circuit size, gemma-2-27B.
 D.3 Linear Latent Faithfulness Detection.
 Here we provide additional information about layer-wise linear probe performance and similarity between faithfulness vectors and with other AI safety linear vectors.
-Figures13, 14, 15, 16, 17, 18, 19, 20, 21 show the layer-wise performance of the faithfulness linear probes for 2-hop reasoning, AGNews and Ledgar.
-As shown in Figure 30, 23 and 31, cosine similarity between task-specific linear faithfulness and AI safety behaviors vectors becomes more pronounced with the size of the model.
+Figures 16-24 show the layer-wise performance of the faithfulness linear probes for 2-hop reasoning, AGNews and Ledgar.
+As shown in Figure 25, 26 and 27, cosine similarity between task-specific linear faithfulness and AI safety behaviors vectors becomes more pronounced with the size of the model.
 Figure 16: Linear faithfulness probe classification performance for 2-hop reasoning, gemma-2-2B.
@@
   • ($C_{1}$) Complete reasoning failure.
-If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\Gamma,o_{2}\notin\tilde{h^{\ell}_{k}}$.
+$\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\Gamma,o_{2}\notin\tilde{h^{\ell}_{k}}$.
 Observation: Wrong prediction, incorrect unfaithful explanation, ground-truth bridge object undetected in circuit $\Gamma$.
@@
   • ($C_{2}$) Internal-external reasoning mismatch.
-If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\exists(k,\ell)\in\Gamma,o_{2}\in\tilde{h^{\ell}_{k}}$.
+$\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\exists(k,\ell)\in\Gamma,o_{2}\in\tilde{h^{\ell}_{k}}$.
 Observation: Wrong prediction, incorrect unfaithful explanation, but ground-truth bridge object detected in circuit $\Gamma$.
-Interpretation: Model appears to have correct internal knowledge but generates inconsistent explanations, indicating potential reasoning-explanation dissociation because of either deceptiveness or hallucination.
+Interpretation: The model appears to have correct internal knowledge but generates inconsistent explanations, indicating potential reasoning-explanation dissociation because of either deceptiveness or hallucination.
   • ($C_{3}$) Explanation-prediction association.
-If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$ and $\widehat{o_{2}}=o_{2}$.
+$\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$ and $\widehat{o_{2}}=o_{2}$.
 Observation: Wrong prediction with unfaithful but correct explanation.
-Interpretation: Model associates correct bridge object with incorrect prediction during explanation generation, suggesting superficial pattern matching without genuinely resolving the first hop of the 2-hop reasoning.
+Interpretation: The model associates the correct bridge object with incorrect prediction during explanation generation, suggesting superficial pattern matching without genuinely resolving the first hop of the 2-hop reasoning.
   • ($C_{4}$) First-hop reasoning failure.
-If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}\neq o_{2}$.
+$\widehat{o_{3}}\neq o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}\neq o_{2}$.
 Observation: Wrong prediction with faithful but incorrect explanation.
-Interpretation: Model consistently follows incorrect reasoning pathway, indicating systematic error in first-hop reasoning or concept misunderstanding.
+Interpretation: The model consistently follows an incorrect reasoning pathway, indicating error in first-hop reasoning or concept misunderstanding.
   • ($C_{5}$) Second-hop reasoning failure.
-If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}=o_{2}$.
+$\widehat{o_{3}}\neq o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}=o_{2}$.
 Observation: Wrong prediction with faithful and correct explanation.
-Interpretation: Model correctly identifies bridge object but fails in second reasoning step, suggesting error occurs after successful first-hop completion.
+Interpretation: The model correctly identifies he bridge object but fails in second reasoning step, suggesting error occurs after successful first-hop completion.
   • ($C_{6}$) Shortcut learning.
-If $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\Gamma,o_{2}\notin\tilde{h^{\ell}_{k}}$.
+$\widehat{o_{3}}=o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\Gamma,o_{2}\notin\tilde{h^{\ell}_{k}}$.
 Observation: Correct prediction with unfaithful incorrect explanation, ground-truth bridge object undetected.
@@
   • ($C_{7}$) Deceptiveness or hallucination.
-If $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\exists(k,\ell)\in\Gamma,o_{2}\in\tilde{h^{\ell}_{k}}$.
+$\widehat{o_{3}}=o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\exists(k,\ell)\in\Gamma,o_{2}\in\tilde{h^{\ell}_{k}}$.
 Observation: Correct prediction with unfaithful incorrect explanation, but ground-truth bridge object detected internally.
-Interpretation: Model possesses correct internal knowledge but generates deceptive (or hallucinated) explanations, suggesting reasoning-explanation dissociation or alternative reasoning pathways.
+Interpretation: The model possesses correct internal knowledge but generates deceptive (or hallucinated) explanations, suggesting reasoning-explanation dissociation or alternative reasoning pathways.
 This case is expected to be rare, otherwise highlighting a case where $f$ is not honest in its self-NLE while ”knowing” the ground truth bridge object, raising a problem in $f$ alignment.
   • ($C_{8}$) Explainer parrot.
-If $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$ and $\widehat{o_{2}}=o_{2}$.
+$\widehat{o_{3}}=o_{3}$, $F(x,e)=0$ and $\widehat{o_{2}}=o_{2}$.
 Observation: Correct prediction with unfaithful but correct explanation.
-Interpretation: Model generates expected explanations without corresponding detectable internal reasoning, suggesting post-hoc explanation generation (i.e.
+Interpretation: The model generates the expected explanation without corresponding detectable internal reasoning, suggesting post-hoc explanation generation (i.e.
 ”explainer parrot” behavior).
   • ($C_{9}$) Alternative reasoning pathway.
-If $\widehat{o_{3}}=o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}\neq o_{2}$.
+$\widehat{o_{3}}=o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}\neq o_{2}$.
 Observation: Correct prediction with faithful but incorrect explanation.
-Interpretation: Model uses consistent but non-canonical reasoning pathways, indicating systematic bias or alternative reasoning mechanism that leads to correct outcomes through unexpected intermediate steps.
+Interpretation: The model uses a consistent but non-canonical reasoning pathways, indicating bias or alternative reasoning mechanism that leads to correct outcomes through unexpected intermediate steps.
   • ($C_{10}$) Reliable oracle.
-If $\widehat{o_{3}}=o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}=o_{2}$.
+$\widehat{o_{3}}=o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}=o_{2}$.
 Observation: Correct prediction with faithful and correct explanation.
@@
 While this taxonomy relies on the interpreter ability to decode the model’s internal activity, the systematic patterns observed across different models and tasks provide convergent evidence for these behavioral categories.
-We give examples of thix taxonomy in Appendix H
+We give examples of this taxonomy in Appendix H
 Incorrect Predictions
@@
 gemma-2-27b 22.8% 4.4% 12.5% 17.1% 43.1%
+mistral-3-7b 46.5% 13.7% 7.1% 8.4% 24.2%
 Table 10: Distribution of categories for incorrect predictions across model sizes.
@@
 gemma-2-27b 12.8% 3.1% 15.3% 6.4% 62.4%
+mistral-3-7b 20.6% 4.0% 19.3% 5.2% 50.9%
 Table 11: Distribution of categories for correct predictions across model sizes.
@@
 Taxonomy Faithfulness Enhancement Analysis.
-Table 9,10 and 11 respectively show the detailed impact of hallucination inhibition, linear faithfulness amplification and deceptiveness inhibition steering on self-NLE faithfulness.
+Table 12,13 and 14 respectively show the detailed impact of hallucination inhibition, linear faithfulness amplification and deceptiveness inhibition steering on self-NLE faithfulness.
 Categories C2 and C7 are the most prone to be turned into faithful self-NLE overall.
@@
 Table 14: Percentage of initially unfaithful explanations that become faithful after deceptiveness steering interventions, by category and model size (third experimental condition).
-As shown in Figures 27, 28, and 29, we observe consistent transition patterns when steering makes unfaithful explanations faithful.
-For initially incorrect predictions, category C1 (complete reasoning failure) consistently transitions to C4 (first-hop reasoning failure), achieving faithfulness approximately 10% of the time under NeuroFaith linear faithfulness amplification and hallucination inhibition.
+As shown in Figures 29, 31, and 33, we observe consistent transition patterns when steering makes unfaithful explanations faithful.
+For incorrect predictions, category C1 (complete reasoning failure) consistently transitions to C4 (first-hop reasoning failure), achieving faithfulness approximately 10% of the time under NeuroFaith linear faithfulness amplification and hallucination inhibition.
 Category C2 (internal-external reasoning mismatch) predominantly transitions to C5 (second-hop reasoning failure) when steering succeeds.
 Category C3 (explanation-prediction association) exclusively leads to C5 (second-hop reasoning failure) upon becoming faithful.
-Initially accurate predictions follow similar patterns.
+Accurate predictions follow similar patterns.
 Category C6 (shortcut learning) transitions to C9 (alternative reasoning pathway) when made faithful, while category C8 (explainer parrot) becomes C10 (reliable oracle).
 Category C7 (deceptiveness or hallucination) can transition to either C9 or C10, though it more commonly becomes a reliable oracle (C10).
-We give several examples of unfaithful self-NLE made faithful in Appendi H.
+We give several examples of unfaithful self-NLE made faithful in Appendix H.
 (a) Initially inaccurate predictions.
@@
 F NeuroFaith Faithfulness Measure Comparison
-In this section we examine how faithfulness as measured by NeuroFaith compares to existing Counterfactual Intervention (CI) Atanasova et al.
-(2023) and Attribution Agreement (AA) approaches Parcalabescu & Frank (2024) for evaluating LLM self-NLE faithfulness.
+In this section we examine how faithfulness as measured by NeuroFaith compares to existing Counterfactual Intervention (CI) (Atanasova et al., 2023) and Attribution Agreement (AA) approaches (Parcalabescu and Frank, 2024) for evaluating LLM self-NLE faithfulness.
 We focus on 2-hop reasoning and begin with a qualitative comparison that illustrates potential differences between these approaches, followed by a quantitative analysis that provides evidence for these observations.
 F.1 2-hop Reasoning Faithfulness Qualitative Comparison.
-To illustrate the critical differences between existing faithfulness evaluation approaches and NeuroFaith, we examine a concrete 2-hop reasoning example (see Figure 32) that illusrrates fundamental limitations in current methods (CI and AA).
+To illustrate the critical differences between existing faithfulness evaluation approaches and NeuroFaith, we examine a concrete 2-hop reasoning example (see Figure 35) that illustrates fundamental limitations in current methods (CI and AA).
 Given input text $x$, prediction $f(x)$, and self-generated explanation $e(x)$, we compare how different faithfulness approaches evaluate the same case.
-The Example Setup:
-  • Original input: $x$ = ”The father of Carol Chomsky is
-  • Original prediction: $f(x)$ = ”Harry Abraham Schatz
+We analyse the following example:
+  • Input: $x$ = ”The father of Carol Chomsky is
+  • Prediction: $f(x)$ = ”Harry Abraham Schatz
   • Counterfactual Intervention: $x_{cf}$ = ”The father of the spouse of Carol Chomsky is
@@
 This evaluation is based solely on the presence of the intervention term ”spouse” within $e(x_{cf})$, establishing consistency between the input modification and explanation content.
-Moreover, Attribution Analysis (AA) methods consists in comparing attribution scores (e.g., using SHAP) to highlight important tokens (e.g.
+Attribution Analysis (AA) methods consists in comparing attribution scores (e.g., using SHAP) to highlight important tokens (e.g.
 ”father”, ”spouse”, ”Carol”, ”Chomsky”) for the prediction and the self-NLE.
 High correlation coefficients between the attribution vectors for prediction and self-NLE would similarly classify the self-NLE as faithful.
-However, NeuroFaith rejects $e(x_{cf})$ as unfaithful because the bridge object (”Morris Halle”) is not contained in the decoded hidden states ($\{\tilde{h^{\ell}_{k}}\}$).
+On the contrary, NeuroFaith rejects $e(x_{cf})$ as unfaithful because the bridge object (”Morris Halle”) is not contained in the decoded hidden states ($\{\tilde{h^{\ell}_{k}}\}$).
 This mismatch between internal computation and self-NLE content reveals unfaithfulness.
@@
 We propose an experimental protocol to evaluate the practical utility of different faithfulness measures for model analysis.
-The protocol is motivated by two common applications of explanations in AI systems: (1) troubleshooting models by identifying the source of errors in wrong predictions Biecek & Samek (2024), and (2) detecting potential biases or shortcuts in correct predictions to ensure they align with expected reasoning processes Ribeiro et al.
-(2016).
+The protocol is motivated by two common applications of explanations in AI systems: (1) troubleshooting models by identifying the source of errors in wrong predictions (Biecek and Samek, 2024), and (2) detecting potential biases or shortcuts in correct predictions to ensure they align with expected reasoning processes (Ribeiro et al., 2016).
 We test whether faithful explanations, as identified by NeuroFaith and CI methods, provide better diagnostic information for these purposes.
@@
 Each category can be further subdivided with respect to faithfulness, enabling comparison between faithful and unfaithful self-NLE within each reasoning pattern.
-In the following we propose 3 metrics based on these 4 categories and assess if faithful self-NLE lead to better model debugging and bias targeting than unfaithful self-NLE.
+In the following we propose 3 evaluation protocols based on these 4 categories and assess if faithful self-NLE lead to better model debugging and bias targeting than unfaithful self-NLE.
 Figure 36: Simplified taxonomy of $f$ behavior in two-hop reasoning, based on the status of the prediction and the self-NLE.
@@
 Finally, the Compound Accuracy Score (CAS) quantifies whether faithful explanations provide better error localization:
-$CAS(A,B,hint1)=log\Bigl(\frac{PR(A,B,hint1,faithful)}{PR(A,B,hint1,unfaithful)}\Bigr)$ (5)
+$CAS(A,B,hint1)=\log\Bigl(\frac{PR(A,B,hint1,faithful)}{PR(A,B,hint1,unfaithful)}\Bigr)$ (5)
 A positive CAS indicates that faithful self-NLE better identify first-hop failures.
+This metric is conceptually close to the In-Context Editing instantiation of the approach proposed by Zaman and Srivastava (2025) to compare faithfulness measures.
+In the following, we extend beyond this proposal with two other metrics.
 Quality Metric gemma-2-2b gemma-2-9b gemma-2-27b
-NeuroFaith CI NeuroFaith CI NeuroFaith CI
-hint1 0.04 -0.07 0.06 -1.10 0.75 -0.99
-hint2 -0.44 -0.55 0.28 -0.03 -0.35 1.25
-$r_{2}\rightarrow r^{\prime}_{2}$ 0.09 -2.62 0.17 -0.32 0.31 -0.26
-Table 12: NeuroFaith comparison to CI (higher is better) across models on 352 CI-compatible samples on 2-hop reasoning.
+NeuroFaith CI AA NeuroFaith CI AA NeuroFaith CI
+hint1 0.04 -0.07 -0.23 0.06 -1.10 0.40 0.75 -0.99
+hint2 -0.44 -0.55 -1.23 0.28 -0.03 0.15 -0.35 1.25
+$r_{2}\rightarrow r^{\prime}_{2}$ 0.09 -2.62 1.32 0.17 -0.32 0.01 0.31 -0.26
+Table 15: NeuroFaith comparison to CI and AA (higher is better) across models on 352 CI-compatible samples on 2-hop reasoning.
 Quality Metric gemma-2-2b gemma-2-9b gemma-2-27b
@@
 Table 16: NeuroFaith evaluation across models on the overall 2-hop reasoning dataset.
-Table 14: Correlation analysis between NeuroFaith and CI.
 Second Hop Hint.
-We now test whether faithful self-NLE better identify second-hop reasoning failures through a targeted intervention.
+We then test whether faithful self-NLE better identify second-hop reasoning failures through a targeted intervention.
 Following the logic introduced above, for each input $x$ having led to a wrong prediction ($\widehat{o}_{3}\neq o_{3}$), we create a modified version $x_{hint2}=(o_{2},r_{2},o_{3},x)$ that explicitly provides the second part of the 2-hop reasoning.
@@
 We test whether faithful self-NLE better identify non-canonical reasoning pathways through the modification of the second step of the reasoning trace ($r_{2}$).
-Intuitively, a non-canonical reasoning pathways is more prone to lead to false reasoning when chaning one step of the 2-hop reasoning.
-For each input $x$ having led to a good prediction ($\widehat{o}_{3}=o_{3}$), we create a modified version $x_{r_{2}\rightarrow r^{\prime}_{2}}=(o_{1},r_{1},{\blacktriangle{}},r_{2}^{\prime},\bullet)$ that change the second relation of the 2-hop reasoning input.
+Intuitively, a non-canonical reasoning pathway is more prone to leading to false reasoning when changing one step of the 2-hop reasoning.
+For each input $x$ having led to a correct prediction ($\widehat{o}_{3}=o_{3}$), we create a modified version $x_{r_{2}\rightarrow r^{\prime}_{2}}=(o_{1},r_{1},\tiny{\blacktriangle{}},r_{2}^{\prime},\bullet)$ that changes the second relation of the 2-hop reasoning input.
 For example:
@@
   • $x_{r_{2}\rightarrow r^{\prime}_{2}}=$ ”The father of the movie maker that directed Persona is”
-If faithful self-NLE accurately reflect internal reasoning, then changing $r_{2}$ should differentially decrease performance for Category C (alternative reasoning pathway) versus Category D (canonical reasoning), and this difference should be stronger for faithful self-NLE.
+If faithful self-NLE accurately reflect internal reasoning, then changing $r_{2}$ should significantly decrease performance for Category C (alternative reasoning pathway) versus Category D (canonical reasoning), and this difference should be stronger for faithful self-NLE.
 We build our third metric following the notations introduced above, based on the Compound Accuracy Score:
@@
 Experimental Results.
-We evaluate our three faithfulness indicators using both NeuroFaith and Counterfactual Intervention (CI).
+We evaluate our three faithfulness indicators to compare NeuroFaith to commonly used Attribution Agreement (AA) and Counterfactual Intervention (CI) methods.
 The Wikidata-2-hop dataset provides natural support for computing $x_{r_{2}\rightarrow r^{\prime}_{2}}$ instances, as it contains multiple reasoning chains involving similar objects.
-Additionally, the dataset includes counterfactual interventions through variations in reasoning chains, enabling straightforward CI computation across multiple instances.
-Our evaluation reveals three key findings.
-First, Table 14 shows that NeuroFaith consistently outperforms CI on the subset of CI-compatible samples.
+Additionally, the dataset includes counterfactual interventions through variations in reasoning chains, enabling straightforward CI computation as in Atanasova et al.
+(2023) across multiple instances.
+We implement AA based on gradient-based attributions as in Wiegreffe et al.
+(2021).
+Due to prohibitive cost for 27-billion models, we only apply AA to Gemma-2-2B and Gemma-2-9B.
+Our evaluation reveals 2 key findings.
+First, Table 16 shows NeuroFaith outperforms CI and AA on average on the subset of CI-compatible samples.
 Second, Table 16 presents NeuroFaith results on the complete dataset, showing positive metric values across all models except gemma-2-2b.
 G Classification Examples
-In this section we give two examples of instances characterized as either faithful or unfaithful in classification case.
+This section gives two examples of instances characterized as either faithful or unfaithful in classification case.
 Figure 37: Example from the AGNews dataset where we detect ”sport event” and ”athletic competition” as relevant concepts from the explanation.
@@
 This bridge object is false and does not appear in the set of natural language interpretations of $f$ latent states.
-However, the ground truth bridge object ”Louis Armstrong” is decoded at token index 11 and layers 13 and 14.
+However, the ground truth bridge object ”Louis Armstrong” is decoded at token index $1$ and layers 13 and 14.
 Figure 40: Example from the Wikidata-2-hop dataset where we observe shortcut learning.